A Bibliography of Publications in Science and Public Policy: 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.03

Title word cross-reference

1990s [Ano92m, Old91].

20th [Mar94b, You91]. 21st [New92].

90s [Rad91].

abandoned [Jev92b]. absorptive [GJ99]. Academia [Sin90]. Academic [BP94, Pac94, Efa93, EKNU92, Hea94, MII90, MI92, Web94b]. Academic-industry [BP94, Pac94, Hea94, Web94b]. Academy [Ano91a, Sin93]. accidental [Mac94]. accomplishment [Duf95d]. account [Gum95a]. accountability [Rap95a, Wei99].

Accounting [CR92a]. Accuracy [Smi93]. achievements [Tri99b]. ACRE [LT93]. activity [Doi94]. Actor [Wey94]. acutely [Edm94]. administrations [Roq95]. advance [Ano92m, Nel91]. Advanced [Bui93, DG92]. Advice [LT93, BC95a, Bra93, Bro93a, Con93, Gum94, Sko93a]. advising [dJ99]. Advisory [Fin99, Joh93, Men99]. aeronautical [Cab92c]. aerospace [HM90]. affecting [Bow95, Sko93b]. Africa [Gro94, Phi90, Tin93, Vit92]. African [For95b]. after [Lüs90]. age [Por93]. Agenda [Met93, Smi94]. Agricultural [Jaf92, Cab94b, GB93, Lac95]. agrofood [BE95]. alignment [KKM99]. alliances [O’D90]. alternatives [Bra93]. Ambitious [Cab92a, Wat94a]. America [BP94].

American [GB93, Jaf92]. among [Roq95]. Analysis [Ric93, BM91, BM92, CM95, Hop99, Jos94, Kar94, LL99, Ran94b, Sch93a, Sim93a, WGBW99, Web94b]. Analytical [Ano91a]. Anamorphosis [Paq90]. ANC [Gro94]. any [Lis95]. Application [Kie95, SP99]. Applied [DN93, Ric93]. appraisal [Mes93]. approach [Ram94a]. approaches [Kos94, TW99]. appropriate [Cab91, Ram94a]. arduous [Baa94]. area [Jos94]. Argentina [Kat91]. arising [Mun95]. Armaments [Blu90]. Arms [Har90]. Asia [Smi94, Car99]. Asian [Liy94]. Aspects [Efa93, Szá94]. aspirations [Kal91]. Asselt [SW95b]. assessed [AC95]. Assessing [Kos94]. assessment [Bec93, Dur99, DKB99, Hen99, Hör99, Sch93b, Von95, Wal92]. associated [Mun95]. assumption [Jev93]. assumptions [Div92c]. ASTEC [Ano90a]. Asymmetric [Div95a]. Athena [EKNU92]. atomic [Lüs90]. attitudes [Zho91]. attraction [Fit93]. Australia [GJ99]. Australian [Ano90a, Joh93, Rya90]. Austria [Ric93]. authors [Ano95h]. aviation [Geo93a]. avoided [Gre92].

B [BC95b, vAR95]. background [Lin90]. Balance [Duf93a, Sir92, Smi91d]. Bank [Sag91]. barracks [Dag93]. Barriers [EKNU92]. based [Acs99, MSdM99]. Basic [Hol92a, PS92]. basis [Phi90]. be [Jev92b]. become [Glo95]. Before [Kne90, Mar90]. behind [Fau95]. Bench [GR99]. Benchmarks [LL95]. benefits [Gre93, Smi91b]. better [May99, Rah90]. between [DKB99, LOFM92, MFGK94, May99, SF95, The95]. beware [Hay92]. Beyond [Vit92, Woo91]. biblical [Mar94b]. bibliography [Pac94]. big [Joh99a]. biodiversity [Jen99a]. Bioinformatics [SB99]. biological [Mey90]. biology [Goo94a]. Biotech [Bab99]. Biotechnology [BE95, Cab94b, WM93, Ber99, Gal94, Isn93, Jaf92, LT91, LT93, Mar94a, Rus92, SF95, Wal93, Web90]. boards [Div92b]. Boehmer [SW95b]. Boehmer-Christiansen [SW95b]. bomb [Coo91, Lüs90]. book [Gre93]. books [End99a]. boundaries [Irw93]. Brazil [Cab92b, EMT98, Mar94a, MS92, MS95]. Brazilian [BdSdOR95, Cab92c, Cab94b, Dag93, Gal94, SF95]. breed [GR99]. Brian [SW95b]. bridge [May99]. Bridging [DKB99, Web94a]. Bringing [Ver94]. Britain [LT93, Twe90]. British [Div91b, Div92a, MIN^{+}90, Utt91]. building [RS94]. business [Lis95]. buy [Ste99].

Call [Irw91a]. Can [Wat94b]. Canada [Isn93, LOFM92, Nio91, dlMP94]. Canadian [Voy99]. capabilities [MSdM99]. capability [KRY94, RS94]. capacity [CB91, GJ99]. capture [Pow95]. Cartographer [Bad90]. case [Acs90, BE95, Fis99, GJ99, Hay93, Myt90, Pow95, Spi92, SB99, Utt91, Wey94]. cent [Hol95a]. Center [Ano91a]. centers [Etz94]. central [Rap95a]. Centre [Ano91b]. century [Mar94b, New92, You91]. CERN [Ric92]. chain [BE95]. challenge [Dec94]. challenges [Ano92m, For95b, Sch93b, WGBW99]. change [BC95b, Eva99, Gor92, How90, Jos99, SW95a, Smi94, vAR95]. changed [The95]. Changes [Jen99a, KT95, Jas91b, Jas94, Szá94]. Changing [Ale92, Kod91, Liy93, For95b, Sko95, Tri99a]. chaos [BL93]. Check [Sap90]. chemical [AK94]. China [Lin90, Yuh92, Zho91]. Chinese [Bao91, Bha90]. choice [Ram94a, Rei90]. choices [CR92b]. Christiansen [SW95b]. cigar [Hol95a]. Circumstantial [dlMP94]. Citation [GWD92, Woo91]. Citizen [Hör99]. citizens [Ber99]. civilian [Har90]. classic [Glo95]. cleaner [You94]. Clearing [Han95]. climate [BC95b, Eva99, SW95a, Smi94, vAR95]. closer [Mol92]. clustering [Voy95a]. Co [Col92, Her94, BE95, Irw92a, Jev93, LOFM92, Ten90]. co-location [Jev93]. Co-operation [Col92, LOFM92, Ten90]. co-ordinated [Irw92a]. Co-ordinating [Her94]. co-ordination [BE95]. Coherent [Geo93a]. cohesion [Mor95]. cold [Duf95a]. collaboration [Blu90, HM90, Hay93]. collaborative [Bui93]. collected [Bud92]. Collection [Ste92, Jen99b, Raj90, Yuh92]. collide [EMT98]. colony [Duf95a]. comes [Mun99]. Comment [ACL94, Smi93]. commentators [Div95b]. Comments [BC95b, SF95, SP99, vAR95]. commerce [Ste99]. commercial [Duc90]. commercialisation [Ogb90, Vis95]. commercialization [GB99]. Committee [Fin99, Wel94]. Common [Muñ90]. communication [Axo99, De 95]. Community [Bar90b, Gon95, LOFM92, Ten90]. companies [PP91, SV91, Sta94]. company [Cab94b]. comparative [Ren95]. comparison [HS91, MII90, MI92, Vis95]. comparisons [Hol91]. Competent [Why94]. Competing [Mah99, Ave93]. competition [Myt90]. competitiveness [AK94, BM91, Duf95c, TW99]. Complex [Koj90, dBtH99]. composition [LT93]. Comprehensive [Sim93a, Sim94a]. Computer [Mac94, OHB92, CM95, Gre93]. Computer-related [Mac94]. Computerisation [Bha90]. computing [Mol90]. concept [JB99]. conceptual [OHB90]. concern [GR99]. concerns [Rap95a]. conferences [AJ99]. conflicting [Sri99]. confused [Ran94a]. consensus [AJ99, Bud90]. conservation [Bun99]. considerations [Bar94a]. Considering [JB99]. consortia [Han95]. constituency [KKM99]. construction [BC95b, SW95a, vAR95]. constructivist [Yea94]. consultation [Fin99]. contacts [Efa93]. contemporary [Div93, Koj90, Nel99a]. context [CB91, Lac95, Lak95]. contexts [Web94b]. contract [Web94a]. Contrasting [Roh93]. Contribution [Gar91, Col92, Hea94]. controlled [OHB92]. Controlling [AC95]. controversies [You91, vE99]. controversy [Mar94a]. conversion [Har90]. cool [Lon93]. coordination [SMMG93]. corporate [RS94]. Correction [Ano92a]. COST [Muñ90]. costs [CR92a, Gre93]. council [Men99, Ano90a]. Countries [Art90, AA91, BP94, Bar94b, Cor94, Jaf92, Kar94, Liy93, MFGK94, RS94, SP99, Woh91, dO91]. country [Duf95a, MHN92]. cover [Bar90a]. cover-up [Bar90a]. crisis [HJD95]. Criteria [Ave93, Wal92]. critic [Mac91]. Critical [Ros91, Gal94, HS91, Irw91a, Rap95b, WM93]. criticism [dJ99]. crossroads [Ric91]. crowding [HS90]. crowding-out [HS90]. crucial [Sim95]. Cuba [Ric94a]. cultural [Fit93, For90, Rah91]. culture [AK94, Gol90, Sch93b, Tin93, Wil90]. cultures [Sin93]. current [WGBW99]. Cyprus [MHN92].

dangerous [Kel92]. dangers [Duf93a]. data [GWD92, Raj90, SWG95]. deaf [vE99]. death [Mac94]. debate [Woo91]. debating [And95]. Decades [Bon91]. Decision [The95, AJ99, Bec93, ES90, JB99, Jos99, May99, dBtH99]. Decision-making [The95, AJ99, Bec93, ES90, JB99, Jos99, dBtH99]. decisions [De 95, Ran94b]. Defence [HS90, Spi92, Buc93, Div92c, EUH90, Hea90]. deficiencies [Bra93]. definitions [DN93]. Definitive [Geo93b]. deliberation [Fis99]. deliberative [Hör99]. Demand [Wal93, ES90, FF93]. democracy [Llo91]. Democratic [Bec93, AJ99, Dar99, Dur99, Fis99, Gro94]. democratize [Ber99]. demonopolisation [dJ99]. departures [MS95]. dependence [EUH90]. dependencies [EMT98]. dependency [HS91]. depolitisation [Dar99]. depth [Wat94a]. Deserves [Glo95]. design [Fau95]. Designing [Joh93]. Destabilising [Gum92]. developed [Woh91]. Developing [Cab91, Cab94a, AA91, Bar94b, Cor94, Liy93, RS94, SP99, dO91]. Development [Ano91a, BM92, DN93, For95b, EUH90, Her94, HS91, How90, Jaf92, Lee91, Liy93, MSdM99, PR90, Sal95, Sig92, Sim95, Smi93, TW99, Wal92]. developments [BL93, SV91]. Dialogues [vE99]. Diary [Ano91c, Ano91d, Ano92b, Ano92c, Ano93a, Ano95a, Ano95b]. did [dlM95]. Dieman [Hol95b]. Differences [Fer95, Kne93, Nel91]. different [End99a]. Diffusion [OHB90, OHB92]. Dilemmas [Jas91a]. dimensional [Sim92]. diplomacy [Eve92]. directions [Von95]. disappointing [Edm94]. Disconnections [SF95]. discovery [Lis95]. discrimination [Hon93]. discussion [Ano90n]. dish [Joh99a]. Disincentives [Pla91]. dislike [CL95]. diversity [Bar90b, Gro90a]. Does [GR99, SB99, vHdP95]. dog [Wel94]. domain [BC95b, SW95a, vAR95]. domestic [Paq90]. dominance [Utt91]. Donald [Smi93]. Dr [Lic91]. draft [Goo94b]. Dutch [BE95, Men99]. dynamics [Fer95, Ric94c].

each [CL95]. East [Ano90b, Ano90c, Ano90d, Ano90e, Ano91e, Ano92d, Ano92e, Ano92f, Ano93b, Ano93c, Bud94, Car99, BP94, Kar94, Mes93, OHB90, Smi90b, Szá94]. East-European [Szá94]. Eastern [Kal91]. EC [BC95a, MFGK94]. Economic [Gre92, Lac95, Lee91, Tri99a]. economics [BL93, Gib93, Sav93]. economies [tP99]. Economy [Ano91a, CH99, Jas91a, Lon94, O’D90, OHB90, OHB92, Rus95, Sch90]. editorial [HJ95]. education [Ver95]. effective [Joh93]. effectiveness [Gov90]. effects [Hol91]. Electronic [Ste99, Fin99]. electronics [BM91]. Elucidating [Hem95]. embraces [Gol90]. EMBRAPA [Cab94b]. emergence [Etz94, Mol94]. emergent [BC95b, SW95a, vAR95]. Emerging [Mol90, RS94, SB99]. empirical [Jon93, Mac94, MFGK94]. end [MS92, Wel94]. end-user [Wel94]. Energy [You91, dO91, Bun99]. Engineering [Ano92l, Bui93, Web90, EKNU92, WM92]. engineers [Div91b]. Enterprises [Spi92, Chu99]. Entertaining [Row94]. entrepreneur [EG99]. environment [LT91, Lon94, Mun95, Sim94b, Sko95, Tri99a, Web90]. environment-friendly [LT91]. Environmental [De 95, Eki92, Eve92, Gib93, Fer95, Gef95, Wat94b]. ephemeral [Jos99]. equipment [HH90]. equity [Bro93b, Kat91]. era [How90]. Error [Ran94b]. ESPRIT [Pow95]. essays [Jen99b, Lic91]. essential [Lon94]. Establishment [Tin93, Hea94, Sin93]. Ethical [Ste91]. Ethics [WM92, For95a, Smi90a]. eulogistic [Hem94]. Eurofighter [SWG95]. Europaea [Sin90]. Europe [AC95, Ano90n, BP94, Bal91, Her94, HJD95, Kal91, Mah99, Mar90, Mol90, Mol92, Muñ90, Sha91, Sin93, WM93, vHdP95]. European [And95, Ano90b, Ano90c, Ano90d, Ano90e, Ano91e, Ano92d, Ano92e, Ano92f, Ano93b, Ano93c, Bar90b, Fer95, FL91, Gon95, Kar94, Lak95, LOFM92, Mol94, OHB90, Ric92, Sin93, Szá94, Ten90, The95]. Evaluating [Hea94, Ave93]. Evaluation [Gov90, WY99, Bar94a, Den94, GWD92, Rap95b, Web94b, WM93]. evidence [Jon93, dlMP94]. Evolution [FL91, EMT98, Gum91, KKM99]. Evolutionary [Goo94a, Szá94, BL93, Dos91, Nel91]. Evolving [Sim94b, Sko95]. excellence [MFGK94]. exchanges [Efa93]. Expanding [Sav93]. expansion [Ing94]. expenditure [Kat91]. expenditures [Hol91].

Experience [Bui93, Ale92, Cab92c, Fin99, Kat91, Kei92, Lee91, Muñ90, Sag91, SF95]. experiences [Sch90, Ten90]. Experimental [DN93, KW94]. Expert [Lev99, Woh91, Gum94]. expertise [Edw99, Gon95, HJD95, Lak95, Ren95, Roq95, RM99, Smi91a, The95, Wei99, dBtH99]. Exploiting [Brz99]. exploration [Mac94]. exploratory [LL99]. export [MFGK94, SV91]. Exposing [Jev92a]. Extending [Fin99, Kar94]. extension [Rap95b]. extra [Hoo93].

F [MII90]. facing [WGBW99]. fact [Tay93]. failed [Div92b]. failure [Div92a]. faith [Vit92]. fallacy [Fry90]. Fallibility [Hud99]. familiarity [GR99]. fantasy [Tay93]. farmers [Bab99]. farming [You94]. Fascinating [Kel92, Geo94]. federal [Atk93, Gef95, Gre93, Isn93]. fields [SB99]. Fills [Buc94]. fin [HJD95]. financed [Sch90]. fine [GB99]. Finland [Ale92]. Finnish [Efa93]. Firm [Doi94, Nel91]. firms [Acs99, Paq90]. first [Mun99]. five [Hol95a]. five-cent [Hol95a]. focus [DKB99]. focused [HS91]. fog [Han95]. Food [Mun99]. force [CR92b]. Forces [Bow95]. foreign [Brz99, EUH90]. foresight [TW99]. Foreword [Bur90]. form [Blu90, Hör99]. formal [CM95]. formulation [Gro94, Sik95, Yuh92]. Forty [Lüs90, Sik90]. Foundation [Gar91, Lin90]. foundations [Pig91]. four [Lüs90]. fragmented [Hem94]. framework [Ren95, Sim92, TW99]. France [MII90, Ric94b]. Freedom [Hay93]. French [GB93, Pig91]. friendly [LT91]. front [Von95]. function [Bec93]. future [Gib94, Sch91]. Fuzzy [Cor95, BM92].

game [Hol95c, Voy95b]. games [Bab99]. gap [Buc94, DKB99, Hay90]. gathering [Gum93]. Gender [Hon93]. Genetic [Fin99, WM92]. genetically [Mun95]. Genome [GR99, Ste91]. German [Kar94, AK94, Her94, Mes93, Sch90, Zac92].

Germany [Con93, MII90]. Gesellschaft [Zac92]. get [End99a]. Getting [Fau95]. giant [Hor94]. giving [Bro93a]. Global [BC95b, Kha91, Rus92, SW95a, Smi94, vAR95, Bar94b, Cab94a, Kod91]. globalisation [How90, RS94]. Globalization [Voy95a]. GMOs [LT91]. go [dlM95]. Going [Gre93]. Goldsmith [Lic91]. good [Hol95b]. governing [Nel99b]. Government [Hay92, Sim95, Bao91, Bra93, Bro93a, CL95, Con93, Eki92, GB99, Irv91, Joh93, Kos94, MII90, MI92, WY99, ES90, Rya90]. government-sponsored [GB99, WY99]. Governments [Div92b, Paq90, Sko93a]. great [Kin91]. greening [LT91]. Group [Hea94, Hon91]. groups [DKB99, Irw92b]. growth [LL95]. Guest [HJ95]. guidelines [Phi90]. Guides [Div95b]. guiding [EES90, Lin90]. Gulf [Wey94].

happened [Ver95]. harmonisation [AC95]. head [Dec94, Eki92]. helicopter [Utt91]. Helix [EL99, SP99]. heritage [For90]. Hero [Str90]. Hi [Cab92b]. Hi-tech [Cab92b]. High [Acs90, DM92, Wey94]. Highly [Gum95a, Bie95, Mah99]. Historical [Div93, Koi93]. history [Hem95, Mil92, Rah91]. honour [Lic91]. hope [Duf93b]. horizons [Sav93]. House [Wil92]. Human [CR92b, GR99, Ste91]. Hungarian [OHB92]. Hungary [Bal92]. Hunt [Che94]. hybridoma [Vis95].

Ibero [BP94]. Ibero-America [BP94]. ideological [Gol91]. IEATP [Bui93]. ignores [Gib94]. II [Mar90]. impact [Bar94a, Den94, Kos94, Mar94a, MIN^{+}90, PR90]. implementation [Sik95]. Implications [Cor94, Cla95, KR93, Kha91, Ste91]. Important [Jos94, Edm94]. Impressive [Sch93a]. Inadequate [Div92c]. income [BP94]. increasing [ES90]. Index [Ano90f, Ano91f, Ano92g, Ano93d, Ano94a, Ano95c, Ano99]. India [Gov90, Rah91, Raj90, Sik95, Vis95]. Indian [Sik90]. indicator [Arc92, Sir92]. indicators [Ano92m, CR92b, DG99, GL92, GWD92, Hol92a, Kei92, NAS92, Yuh92, dlM92]. Individual [Cap90]. Industrial [BM91, Kat91, Bow95, CM95, Etz94, EG99, Kod91, LL95, Voy95a, WY99]. industries [AK94, Doi94]. Industry [AK94, And90, BP94, Bha90, BM91, Cab92c, CL95, Fau95, Hea94, Jas94, Kne99, Pac94, Web94b, WGBW99]. industry-public [Fau95]. ineffectiveness [Isn93]. Influence [Sri99, Gef95, Hol91, Wat94b]. informatics [MS92]. Information [MHN92, Geo94, TL99, Why94, Bui93]. informative [Bie95]. informing [Ber99]. infrastructure [MSdM99]. Initiative [Mol92, Mol94]. Innovation [Bud92, Arc92, Art90, BM92, Car99, Chu99, DG99, EES90, EMT98, Gef95, GJ99, Irv91, Jas91a, Jev93, KR93, Kei92, LL95, Nel99a, Nio91, tP99, Pav92, Rya90, SP99, Smi91d, Sun95, Wal93, Zim91]. Innovations [Phi90, OHB90]. innovative [RS94]. inquiry [Fis99]. insight [Tri99b]. Inspiration [Div91b]. Institute [Ric93]. institution [Ric91]. Institutional [Web90]. Institutionalised [dJ99]. institutions [Sch90, Web94a]. instruments [Irv91]. Integrated [DKB99, Ram94a]. Intellectual [KRY94, KI93, Kne99, Ram94b]. Intelligence [Gum93]. intelligent [TL99]. Inter [Dre90]. Interactive [Cla95]. interchange [SWG95]. interconnections [Gro90b]. Interesting [Jen99b]. interests [Sri99]. intermediary [Edw99]. internalisation [Pow95]. International [Gar91, Gro90b, HM90, Hay93, Hea94, Ten90, WGBW99, Web94b, And90, Art90, Col92, DG99, Hol91, Hor94, MI92, Myt90, Rus95, Sko93b, Sko95, Vis95, Hay93, Ric93]. Internationalisation [Efa93, Pav92, ES90].

Internationalising [Sig92]. Internationalist [May91]. Internationalization [Paq90, SdlMK90]. Internet [Fin99]. interpretation [Kne93]. Intervening [Smi93]. interventionist [Etz94]. Introduction [Hea90, Sin90, Lic91]. Inventing [Smi93]. Invention [Edg91]. investment [Jaf92, MSdM99]. island [TL99]. issue [Hea90]. Issues [Ing94, Baa94, Gre92, Raj90, Ros91, Ste91].

Jan [SW95b]. Janeiro [EMT98]. Japan [Han95, Kne99, Kod91, MII90, Mol90, Roh93, Sig92, WGBW99]. Japanese [Bar95, BM91, BM92, Div92a, Doi94, Koi93]. joint [Koi93]. journalism [Irw91a]. justice [JB99].

key [Baa94]. Knowledge

[Bar94b, For95a, Cla95, CH99, Div93, MSdM99, May99, Rus95, Ver94, Woh91]. knowledge-based [MSdM99]. Korea [Lee91]. Korean [Chu99].

laboratories [CB91, Tri99a]. laboratory [KW94, Ric92]. labs [Dag93]. Lack [Irw92a, Jon93, Mor95, Ram94b]. Lacking [Gum94, Wat94a]. Land [Hol95b]. large [Bal91, FF93, Mol94]. large-scale [FF93, Mol94]. largest [PP91]. lasting [Jos99]. Latin [Jaf92]. Latvian [KT95]. Laws [AA91]. Laying [Pig91]. lead [SB99]. leadership [SMMG93]. leap [Vit92]. Learning [Car99, Atk93]. led [Twe90]. legacy [Kne93]. less [Woh91]. lesson [Car95]. Lessons [Her94, Atk93, Irv91, Sag91]. letters [SW95b]. Letting [Wel94]. level [And95]. liability [Mun95]. Liberty [Mey90]. Lichfield [Lic91]. life [Fre91, Nob93]. light [Div91a]. linkage [Fau95]. Linking [Mar93]. literature [Ano90b, Ano90c, Ano90d, Ano90e, Ano91e, Ano92d, Ano92e, Ano92f, Ano93b, Ano93c, Bud92]. little [Tri99b]. livestock [Cab94b]. location [Jev93]. logic [BM92, Nob93]. London [Hon91]. Look [Bon90, LL95]. Lords [Wil92]. Lucky [Duf95a].

machinery [OHB92]. MacKenzie [Smi93]. main [Ric94b]. major [Buc94]. maker [Hol95a, Rad91]. making [AJ99, Atk93, Bar94a, Bec93, Edw99, ES90, GWD92, Hop99, JB99, Jos99, May99, Men99, Met93, Smi94, Sri99, The95, WGBW99, dBtH99]. Man [Lüs91, Duf93b]. Management [BdSdOR95, Bal92, FF93]. Manchester [Ano92l]. manipulated [Mun95]. manufacturing [Doi94, DG92]. map [Gol91]. Mapping [GR99]. Marjolein [SW95b]. market [Jas91a, SF95, Utt91]. Marketing [dSD95]. markets [Wal93]. Maryland [Acs90]. mass [HS91]. masses [Ver94]. materials [Kem93]. maths [Bon90]. matter [vHdP95]. Maurice [Lic91, Lic91]. Max [Zac92]. Max-Planck-Gesellschaft [Zac92]. Maximising [Smi91b]. measure [Hol92b]. measurement [Woo91]. mechanisms [PS92, WGBW99]. Medical [Smi90a]. Medicine [Smi91c, Smi90b]. medicines [AC95]. medium [Chu99, Woh91]. medium-sized [Chu99]. Meetings [Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano91g, Ano91h, Ano91i, Ano91j, Ano92h, Ano92i, Ano92j, Ano92k, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano95d, Ano95e, Ano95f, Ano95g]. meets [Smi90b]. megaprojects [Ave93]. Merging [Rus95]. merits [Duf93a]. Meske [Kar94]. Methodological [Bar94a]. Methodology [Raj90]. methods [CM95]. meticulously [Sim94a]. Mexico [CL95]. Microprocessor [Mol92]. middle [BP94]. middle-income [BP94]. mighty [Ric92]. migration [KI93]. military [Har90, Smi93, Dag93]. military-civilian [Har90]. milk [Hor94].

Ministry [WGBW99]. minor [Glo95]. missed [SB99]. mobilising [Sal95]. Model [Eva99, Cro94, Dur99, MFGK94, Zim91]. models [Jev93]. modern [For95a]. modernity [Hen99]. modes [Myt90]. Mongolia [Baa94]. mother [Bro93b]. moves [Mol92]. moving [TL99]. muddled [Jev92a]. Museums [Fit93]. must [Jev92b]. mutual [BC95b, SW95a, vAR95]. Myth [Gum95b]. myths [Wil90].

nano [Por93]. nano-age [Por93]. Nanotechnology [PR90]. Nathaniel [Lic91]. National [Tri99a, DG99, GJ99, Nio91, tP99, Lin90]. Natural [Lin90]. nature [Cla95, dlM92]. necessity [Bro93b, Mey90]. needed [Jen99a]. neglected [Jos94]. net [Zim91]. Netherlands [MII90]. Network [Liy94, Art90]. networks [Acs90, EES90, Kie95, Sim93b, Wey94]. neural [Kie95, Mol90, Zim91]. Nigeria [Ogb90]. No [Bud90, Bon90, Irw93]. non [Mar92]. non-violent [Mar92]. Norway [Ano91b]. Norwegian [Sko93a]. note [dlMP94]. Notes [Ano95h]. notions [Rap95a]. novel [Fau95]. NTBFs [Acs99]. nuclear [Wel94]. numerically [OHB92].

objectives [Cab92a]. observations

[OHB90]. OECD [Ano92m, Hol95c, tP99]. off [Sta94]. OMI [Mol92]. One [Gol90, New92]. only [Hol95c]. Open [Mol92]. opening [Por93]. opens [Lic91]. operating [PS92]. operation [Col92, LOFM92, Ten90]. opinion [Roq95]. Opportunities [MSdM99, SB99]. options [Gov90]. order [Kod91, Wal91]. ordinated [Irw92a]. ordinating [Her94]. ordination [BE95]. organic [You94]. organisational [Bar94a]. organisations [Web94a]. organisms [Mun95]. orienting [Ric94b]. original [Cab95]. other [CL95, Kar94]. our [Gol91]. output [MIN^{+}90]. Overburdened [Mor94].

palaver [Paq90]. panels [Hör99]. paper [BC95b, vAR95]. paradigm [Rui94]. paradigms [Now94, Rus95, Sun95]. paradoxes [RM99, Wei99]. parliament [Llo91]. participation [Jos99]. Participatory [Dur99, Hen99, Fis99]. particle [Ric92]. partnering [Myt90]. Past [Tri99b, Div94c]. Patent [GB99]. Patenting [Arc92, Mar94a]. Path [HS91, EMT98]. Path-dependency [HS91]. paths [Car99]. patterns [RS94]. payments [Sir92]. peaceful [Sch91]. People [Zho91]. perception [Nob93]. perestroika [Kne90]. performance [Hol92a, MFGK94]. period [Jas94]. perspective [Bar94b, Gum91, Kod91, Mah99, O’D90, Sag91, Sko93a]. Perspectives [Dos91, Ing94, Liy93]. pharmaceuticals [Mar94a, Smi91d]. Phase [Rui94]. Phases [Now94]. phenomenon [Jos99]. physicist [Lüs90]. physics [Ric92]. place [GJ99]. places [LL95]. Planck [Zac92]. planet [Mot91]. planning [NAS92]. plant [Rap95b]. Playfulness [Irw91c]. Playing [Lon93]. Plugging [Hay90]. poet [May91]. points [Koi93]. Poland [Jas91a]. policies [Acs99, Cab91, Chu99, Eva99, GB99, Pig91, SF95, Yuh92]. Policy [Ano90m, Ano92l, Gar91, Hon91, Hop99, Liy94, May99, TW99, AK94, Ano90n, Ano91k, Atk93, Baa94, Bar94a, Bar95, Bec93, BC95b, BC95a, Bro93b, Cab92c, CR92a, CR92b, CM95, Con93, Cro94, Dag93, De 95, Div93, Div95b, DKB99, Edw99, Etz94, EMT98, EG99, Fer95, FL91, For95b, Gal94, GWD92, Gef95, Geo93a, Goo94a, Gov90, Gro94, Gum91, Hol95a, Ing94, Irv91, Irw92a, Isn93, Jas91a, Jas94, Joh93, JB99, Jos99, KR93, KRY94, Koi93, Lac95, Lak95, Lee91, MSdM99, Men99, Met93, MS92, MS95, Muñ90, Now94, Phi90, Pow95, Rad91, Ran94b, Ros91, Rui94, RM99, Sag91, SMMG93, Sch90, Sch93b, SW95a, Sik95, Sim95, Sin93, Sko95, Smi94, Sri99, Von95, Voy99, WGBW99, Wat94b, WM93, Wil90, Woo91, dO91, dlMP94]. policy [vAR95, vE99]. policy- [JB99, Jos99]. policy-maker [Hol95a]. policy-making [Atk93, Edw99, GWD92, Men99, Smi94, Sri99, WGBW99]. Polish [Jas91b, Jas94]. Political [Kri92, For95b, Roq95, Rus95, Wei99, dlM92]. politics [Bel99c, Hop99, Smi94, The95, Wei99, You91]. Pollution [Irw93, BC95a]. poor [Bab99]. popular [Gum95a]. Post [Yea94]. Post-constructivist [Yea94]. potato [BE95]. Potential [HH90, MHN92, Smi91b]. Potentially [Edm94]. Powell [Wel94]. Power [Smi91a, Zap95, Bab99, BdSdOR95, Div95a, Hop99]. powers [Roq95]. practice [Bro93a, MHN92, WY99]. Preparing [Ogb90]. present [Lin90, Tri99b]. presented [Joh99a]. preserve [Jen99a]. pressure [Irw92b]. price [Lis95]. principles [EES90, Lin90]. priorities [Bud90]. Private [Slo91, Rap95a]. private-sector [Rap95a]. Problems [Ano91a, Buc93, De 95, Gib94, Kri92, Raj90]. procedural [Gre93, JB99]. process [BM92, Den94, Gro94, OHB90]. processes [Gon95, Nel99a, dBtH99]. procurement [Gre93]. production [Cab94b, Utt91]. products [DM92, LT91]. Programme [Bui93, WY99, Ano92l, Art90]. programmes [Dag93, Den94]. progress [Duf95b, Fre91]. Project [GR99, SWG95, Ste91]. projects [Bal91, FF93]. Prometheus [Str90]. Promise [Div94b]. Promoting [Irv91, WY99]. property [KRY94, Kne99]. proposal [CM95]. proposals [Kha91]. prospects [Kei92, Wil90]. protection [KRY94, Rap95b]. provisional [Mes93]. provocation [Irw91c]. provoking [Rep90, Row94]. Public [Acs99, EG99, FF93, Jos99, AK94, Dur99, Edw99, Fau95, Fin99, Gov90, Jas91a, JB99, Kat91, Lac95, Nob93, Rap95a, Roq95, RM99, Slo91, Smi94, Wal93, Wil90, Zho91]. public- [Rap95a].

quality [Fre91]. Quantifying [Ano92m]. quantitative [GWD92, Kos94]. Quebec [GL92]. quest [Sal95]. question [Axo99]. questions [Mar94a]. Quo [Ric91].

R [MII90]. R&D [Ave93, Bar94a, Buc93, CB91, Doi94, ES90, Han95, HS90, Hea90, Hol91, IS94, Jas91b, Kar94, May99, Myt90, tP99, RS94, SdlMK90, Ten90, Vis95, Von95, WGBW99, Wil92]. race [Voy95b, vHdP95]. Radical [Gef95]. radioactive [Kem93]. rationalisation [Den94]. Re [Ric94b]. Re-orienting [Ric94b]. readable [Gum95a]. readings [Cor95]. realities [Wil90]. reality [Kal91, SF95, TL99]. Reawakening [For90]. reconsidered [Str90]. Reconstructing [Goo93]. Reflections [BC95a, Con93, Div94c]. regional [Voy95a]. regulation [LT93]. regulations [AA91]. regulator [Irw91b]. regulatory [AC95, Gon95]. related [Mac94, MII90, MI92]. Relation [MFGK94, The95]. relations [BP94, Div95a, Hea94, Pac94, Web94b]. relative [Isn93]. relay [Voy95b]. releases [Mun95]. relevant [DG99, Mar94a]. remit [LT93]. renaissance [Sha91]. required [Geo93a]. requires [Hoo93]. Research [Ano90m, Ano92l, DN93, Ric94a, Wal92, Bow95, Brz99, Bui93, Cab94b, CR92a, CR92b, EUH90, Fau95, GB99, Her94, HH90, HS91, How90, Jaf92, Koi93, Kos94, KT95, KW94, MII90, MI92, Mes93, Ogb90, Pla91, Rap95a, Ric94b, Ros91, Sag91, Sch90, Sig92, Smi90a, TW99, Vit92, Web94a, Wel94]. researched [Goo94b, Sim94a]. Reshaping [Gol91]. resolving [RM99]. resources [Bar95, CR92b]. respected [Wil93]. respecter [Irw93]. response [Hen99, Now94, SW95b]. responsibilities [Ale92, Dre90, Koj90]. responsibility [Ano90n, Cap90, Lüs90, Rap99]. restated [Kin91]. Restructuring [IS94, Den94, Kat91, Mes93]. results [Ogb90]. rethink [Bro99]. Rethinking [Nas95, CR92a]. retrospective [Kos94]. revealed [Kri92]. review [Arc92, DG99, Gal94, Hol95b]. revised [DN93]. Rewards [Smi91d]. Rich [Bie95, Sim93b]. Richness [Gro90a, Nel99a]. rights [Kne99]. rigour [Ram94b]. Rio [EMT98]. rise [Gol92, KKM99]. Risk [Mun95, De 95, FF93, Wey94]. Risks [Dar99, KW94]. role [Bui93, CH99, De 95, Edw99, ES90, Nel91, Web94a, Yuh92]. Romania [IS94]. Rotmans [SW95b]. Rough [Goo94b]. RTD [Den94]. Rugby [Voy95b]. rules [Ave93, Gre93].

S&T [Ano91k, Ano92m, Cab91, Dag93, Duf93a, Ste92, Yuh92]. S. [BC95b, vAR95]. safe [Kem93]. Saharan [Vit92]. Saving [Mot91]. scale [FF93, Hol91, Mol94, The95]. Scenario [AJ99]. Scenarios [PR90]. scholar [Cap90]. scholarly [Efa93]. Science [Ano90a, Ano90m, Ano91a, Ano92l, Bra93, Cab92c, Cro94, Gar91, Hon91, Jas94, Lin90, Liy94, Llo91, Mar90, Mar92, Rah91, Raj90, Rap99, Sag91, Sik95, Sko93a, Wil90, And95, Ano90n, ACL94, Baa94, Bal92, Bao91, Bar95, Bar94b, Bel99a, BC95b, Bro93a, CM95, Col92, Con93, Dur99, DKB99, End99b, ES90, EKNU92, EG99, FL91, GL92, GB93, GWD92, Goo93, Gro94, Gum91, Hol92b, Hop99, Hud99, Irw91a, Joh93, JB99, Jos99, LL99, Liy93, Mac91, MIN^{+}90, Nob93, Now94, Pig91, PS92, Ric94c, Ros91, Rui94, Sal95, SMMG93, SW95a, Sik90, Sin93, Sko93b, Sko95, Smi91c, Smi94, The95, Tin93, Von95, Voy99, Wat94b, Wei99, Woo91, Yea94, Zho91, dlM92, dlMP94, vAR95, vE99]. Science-and-society [Mar90]. science-policy [BC95b, SW95a, vAR95]. Science-Technology [Ano91a]. sciences [De 95, Lac95, Ano91a]. Scientific [Bar90a, Edw99, Gon95, LOFM92, Lis95, Roq95, RM99, Wei99, dBtH99, BC95a, Efa93, Hol92a, HJD95, Irv91, Kne93, Lak95, MFGK94, Muñ90, Ogb90, Ren95, Ver94, dJ99]. scientist [Dre90]. scientists [GR99, Koj90]. scope [RS94]. Scotland [KKM99]. Seamed [Mar95]. seamless [Mar95]. search [DG99]. Secret [Bel99a]. sector [BdSdOR95, Fau95, Rap95a]. security [Col92]. selected [Pac94]. Selectivity [Rya90]. Selling [Bun99]. sense [Duf95b, Hop99]. Services [DG99, Nas95]. set [Cor95, Lic91]. setting [Smi94]. Shackley [BC95b, vAR95, SW95b]. shared [CL95]. sharing [HH90]. shark [Ano91k]. Shifting [Rap95a]. siècle [HJD95]. Simon [SW95b]. Singapore [TL99]. six [Sap90, Art90]. size [Doi94]. sized [Chu99]. skilled [Mah99]. skimmed [Hor94]. SKY [Gol92]. small [Chu99, MHN92, O’D90]. SME [O’D90]. Snows [dlM94]. Social [Ano90n, Ale92, Bro93b, De 95, Div94b, Goo94a, Kat91, Kri92, KW94, Ste91]. Society [Ano91b, KW94, Fis99, For95b, Lis95, Mar90, New92, Nob93, Rah90]. Socio [Ano91a, Lac95]. Socio-economic [Lac95]. Socio-Economy [Ano91a]. sociologists [Pal94]. sociology [Woo91]. Sociotechnical [KKM99]. solar [Bal91]. Some [CR92a, DN93, Irv91, Rap95a, Sch90, Cla95, Jaf92, Kne93, Mar94a, OHB90, Sag91]. something [Hoo93]. Sonja [SW95b]. soul [Bad90]. South [Gro94, Phi90]. Soviet [Kne93, PS92]. Space [Rei90, Hay93, Lüs91, Zap95]. Spanish [ACL94, SV91, SMMG93]. speaking [Hop99]. special [Hea90]. spending [MII90, MI92]. sphere [Edw99]. Spin [Sta94]. Spin-off [Sta94]. sponsored [GB99, WY99]. SPP [Ano91f, Ano93d, Ano94a, Ano95c, Ano99].

SPRU [Old91]. stability [Web90]. Stable [Lon94]. Stages [Lee91]. state [Etz94, Sch90, EMT98]. state-financed [Sch90]. states [Atk93, Cro94, EG99, LL99]. Station [Hay93]. statistical [Raj90, Sch93a]. statistics [Buc93, Mor94, Ste92, Wil92, Yuh92]. status [Vis95]. STEPAN [Liy94]. still [vHdP95]. Stimulating [WGBW99]. stimulation [Bon91]. Strategic [O’D90, ES90, Myt90, NAS92]. strategies [Lac95, PP91]. strategy [And90, Mol92, Ric94b]. streams [KR93]. Strengthening [Bar94a]. Striking [Smi91d]. Structure [CB91, tP99, Rus95]. structures [For95b, PS92]. struggle [Mar92]. studies [BL93, Yea94]. Study [Hea94, Acs90, BE95, Geo93b, Spi92, SB99, Utt91]. Style [Ren95]. sub [Vit92]. sub-Saharan [Vit92]. Subterranean [Tay93]. success [Div92a, Isn93]. Successful [Cab95, Atk93]. Superficial [Ran94a]. Support [Hon91, Acs99, Bao91, Rya90]. Surplus [Hor94]. survey [GL92]. Surveys [DG92, DG99]. Sussex [Ano90m]. Sustainability [Hoo93]. Sweden [Ing94]. Swedish [And90]. synergy [Har90]. system [Bal92, Bal91, Her94, IS94, Jas91b, Joh93, KT95, Mes93, Nio91, Raj90, SB99]. Systems [Mol92, Ric93, Cla95, GJ99, tP99, Woh91].

Tackling [Dec94]. tail [Ano91k, Wel94]. Taiwan [WY99]. tanks [Men99]. tapestry [Sim93b]. targets [Dar99]. tasks [Bar94b]. tastefully [Joh99a]. taxonomy [tP99]. teaching [Goo93]. tech [Cab92b]. Technical [SWG95, Hen99, Lak95, Nel91, Smi91a, Ver95]. TechnoBodies [Bel99b]. Technological [Fis99, Kei92, PP91, Sir92, Ano92m, Arc92, BM92, FF93, Har90, Her94, Hol92a, KR93, MFGK94, Pav92, Phi90, Rus95, Sch93b, Smi93]. technologies [Gef95, Gum92, Wey94, Woo91]. Technology [Ano90a, Ano91a, Ano91b, Bui93, Duf95c, Etz94, Fre91, Liy94, NAS92, Rad91, Sch90, Spi92, Acs90, Acs99, AA91, And95, Atk93, ACL94, Baa94, Bal92, Bao91, Bar95, Bar94b, Bar90b, Bec93, BdSdOR95, BL93, BM92, Bra93, Bro93b, Cab92c, Col92, Con93, Cor94, Cro94, Div94b, Div95b, DM92, DG92, Dur99, EUH90, ES90, EG99, FL91, For95b, GWD92, Gro94, Gum91, Hem95, Hen99, HS91, Hör99, Jas94, Joh93, JB99, Jos99, KRY94, Kne99, Lee91, Liy93, MHN92, MS95, Mol94, Raj90, Ram94a, Sag91, Sal95, SV91, SMMG93, Sch93b, Sir92, Sko93a, Sko93b, Sko95, TL99, Utt91, Vis95, Von95, WGBW99, Web94a, Wel94, Zho91, dO91, vHdP95, Ano92l]. technology-based [Acs99]. technology-policy [Div95b]. telemedicine [KKM99]. Ten [Atk93]. tene [Dre90]. tenure [Duf95d]. test [MFGK94]. Testing [Fin99]. their [GWD92, Woo91]. theoretical [MFGK94]. theory [BL93, Bro93a, Dos91, Gum94, Nel91, Sun95]. there [Div91a]. Think [Men99]. Thinking [And95, Jev92a]. Third [Cla95, GB93]. Thirty [Voy99]. Thought [Rep90, Row94]. thought-provoking [Row94]. threat [Nob93]. Three [End99a, Sim92, Sun95]. Three-dimensional [Sim92]. Time [Gor92, Bro99]. timely [Sim93a]. times [Mar94b]. tips [Rad91]. today [Ric94a]. together [Hop99]. Too [Hem94]. tool [Pow95]. town [Hol95c]. toxicology [AC95]. Trade [DM92, WGBW99]. trading [Cro94]. tradition [Edg91]. training [Div92b]. trajectory [EG99]. transboundary [BC95a]. transfer [AA91, Bar94b, Cor94, Kne99, Sch90, Sir92, Spi92, Utt91, Web94a, Woh91]. transformation [Rah91]. Transforming [PS92]. Transition [Bal92, Baa94, Jas91a, Jas94, Kin91]. transport [Kem93, Mar93]. trasport [Eva99]. Trends [Cor94, MII90, MI92, LL99, MIN^{+}90, Sko93b, Wal92]. Triad [Wal91]. tribes [Joh99b]. Triple [EL99, SP99]. Trondheim [Ano91b]. troubled [Irw91b]. Truly [And90]. Truth [Jev92b, Hop99]. trying [End99a]. tune [GB99]. turning [Koi93]. Two [Joh99b, KR93, Dre90, Gol90].

UK [Ano92l, Buc93, Bui93, EUH90, Fin99, Gum91, MII90, Mol90, Rap95a, Ste91, Wel94]. UK/Europe [Mol90]. Ukraine [KI93]. unbound [EKNU92]. uncertain [Sal95, Sim94b]. unconvincing [Rep90]. uncovered [Gum95b]. under-researched [Goo94b]. Understanding [Mol94, Dur99]. understandings [Lev99]. Unesco [Ric91, Ric94c]. unification [Zac92]. uninitiated [Div94a]. Uninventing [Coo91]. Union [Fer95]. Unit [Ano90m]. United [Cro94, EG99, LL99]. universities [Sta94]. University [Ano91b, Ano92l, CL95, Ale92, CH99, GB99, Kne99, Ano90m]. university-industry [Kne99]. untouchable [Wil93]. upgrading [WY99]. uptake [CM95]. USA [CB91, Etz94, Kod91, MII90, Mol90]. use [GWD92, Woo91]. user [Che94, Wel94]. users [Bui93, Mol92]. using [Ren95]. USSR [Ano91a]. Utilisation [Lak95]. Utility [Vit92, Nas95]. Utopian [Fry90]. utrumque [Dre90].

vadis [Ric91]. values [Nob93]. Vans [Duc90]. variety [Rap95b]. vehicles [Duc90]. vertical [BE95]. via [Fin99]. vicissitudes [SMMG93]. view [Cab94a, KI93, Roh93, Von95]. Viewpoint [Jas94]. villain [Str90]. violent [Mar92]. Virtual [Bel99c]. Vision [Gib94, Rah90, New92, TL99, WGBW99]. volume [Ano90f, Ano92g]. vs [Slo91].

wag [Wel94]. Wall [Bud94]. War [Mar90, Wey94]. warfare [Rap99]. warming [Kha91]. was [Div91a]. way [Twe90, dlM95]. Wealth [Geo94]. Weapon [EES90]. weaving [Why94]. Web [Bar95]. webs [Mar95]. West [Smi90b]. wheat [Cab94b]. Where [Irw92b, Pal94]. Which [dlM95]. Who [Nel99b]. Whose [EL99]. Wide [Bar95]. Will [Ber99]. Wissenschaft [New92]. within [OHB92]. Women [LL99, EKNU92]. work [CR92b, Wil92]. workshops [AJ99]. World [Bar95, Sag91, Wal91, For95a, Koj90, PP91, Ric94c, Cla95, GB93, Mar90]. would [Ste99]. Wynne [BC95b, vAR95, SW95b].

yardsticks [LL95]. years [Lüs90, Sik90, Voy99]. yesteryear [dlM94].

References

Al-Ali:1991:LRT

[AA91] Salahaldeen Al-Ali. Laws and regulations on technology transfer to developing countries. Science and Public Policy, 18(5): 295–300, October 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Abraham:1995:CME

[AC95] John Abraham and Michael Charlton. Controlling medicines in Europe: The harmonisation of regulatory toxicology assessed. Science and Public Policy, 22(6):354–362, December 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Avila:1994:CSS

[ACL94] J. Avila, M. Candela, and V. Larraga. Comment on Spanish science and technology. Science and Public Policy, 21(3): 184, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Acs:1990:HTN

[Acs90] Dr Zoltan J. Acs. High technology networks in Maryland: a case study. Science and Public Policy, 17(5):315–325, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Acs:1999:PPS

[Acs99] Zoltan J. Acs. Public policies to support new technologybased firms (NTBFs). Science and Public Policy, 26(4):247– 257, August 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Andersen:1999:SWC

[AJ99] Ida-Elisabeth Andersen and Birgit Jæger. Scenario workshops and consensus conferences: Towards more democratic decisionmaking. Science and Public Policy, 26(5):331–340, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Adams:1994:ICP

[AK94] Richard C. Adams and Don E. Kash. Industry culture, public policy, and competitiveness: the US and German chemical industries. Science and Public Policy, 21(5):309–320, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Alestalo:1992:CSR

[Ale92] Dr Marja Alestalo. Changing social responsibilities of the university: The experience of Finland. Science and Public Policy, 19(1):43–54, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Andersson:1990:TIS

[And90] Curt Andersson. Truly international: a strategy for Swedish industry. Science and Public Policy, 17(5):311–314, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Andre:1995:TDA

[And95] Michel André. Thinking and debating about science and technology at European level. Science and Public Policy, 22(3): 205–207, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:AAS

[Ano90a] Anonymous. ASTEC: The Australian Science and Technology Council. Science and Public Policy, 17(6):391–392, December 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:EELa

[Ano90b] Anonymous. East European literature. Science and Public Policy, 17(2):126–128, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:EELb

[Ano90c] Anonymous. East European literature. Science and Public Policy, 17(3):205–206, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:EELc

[Ano90d] Anonymous. East European literature. Science and Public Policy, 17(4):261–263, August 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:EELd

[Ano90e] Anonymous. East European literature. Science and Public Policy, 17(6):393–394, December 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:IV

[Ano90f] Anonymous. Index to volume 17, 1990. Science and Public Policy, 17(6):395–396, December 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:Ma

[Ano90g] Anonymous. Meetings. Science and Public Policy, 17(1): CO3, February 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:Mb

[Ano90h] Anonymous. Meetings. Science and Public Policy, 17(2): co3, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:Mc

[Ano90i] Anonymous. Meetings. Science and Public Policy, 17(3): co3, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:Md

[Ano90j] Anonymous. Meetings. Science and Public Policy, 17(4): CO3, August 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:Me

[Ano90k] Anonymous. Meetings. Science and Public Policy, 17(5): CO3, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:Mf

[Ano90l] Anonymous. Meetings. Science and Public Policy, 17(6): CO3, December 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:SPR

[Ano90m] Anonymous. Science policy research unit at sussex University. Science and Public Policy, 17(3): 207–208, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1990:SRS

[Ano90n] Anonymous. Social responsibility and science policy in Europe: a discussion. Science and Public Policy, 17(2):88–95, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1991:ACP

[Ano91a] Anonymous. Analytical Center for Problems of Socio-Economy and Science-Technology Development, USSR Academy of Sciences. Science and Public Policy, 18(4):263–264, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1991:CTS

[Ano91b] Anonymous. Centre for Technology and Society, University of Trondheim, Norway. Science and Public Policy, 18(3): 197–198, June 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1991:Da

[Ano91c] Anonymous. Diary. Science and Public Policy, 18(5):CO3, October 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1991:Db

[Ano91d] Anonymous. Diary. Science and Public Policy, 18(6):CO3, December 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1991:EEL

[Ano91e] Anonymous. East European literature. Science and Public Policy, 18(5):331–332, October 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1991:IS

[Ano91f] Anonymous. Index, SPP 1991.

Science and Public Policy, 18(6): 427–428, December 1991. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1991:Ma

[Ano91g] Anonymous. Meetings. Science and Public Policy, 18(1): CO3, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1991:Mb

[Ano91h] Anonymous. Meetings. Science and Public Policy, 18(2): CO3, April 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1991:Mc

[Ano91i] Anonymous. Meetings. Science and Public Policy, 18(3): CO3, June 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1991:Md

[Ano91j] Anonymous. Meetings. Science and Public Policy, 18(4): CO3, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1991:STS

[Ano91k] Anonymous. The shark’s tail and S&T policy. Science and Public Policy, 18(5):270, October 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:C

[Ano92a] Anonymous. Correction. Science and Public Policy, 19(2): 130, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:Da

[Ano92b] Anonymous. Diary. Science and Public Policy, 19(2):CO3, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:Db

[Ano92c] Anonymous. Diary. Science and Public Policy, 19(6):CO3, December 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:EELa

[Ano92d] Anonymous. East European literature. Science and Public Policy, 19(1):67–68, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:EELb

[Ano92e] Anonymous. East European literature. Science and Public Policy, 19(2):131–132, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:EELc

[Ano92f] Anonymous. East European literature. Science and Public Policy, 19(3):195–196, June 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:IV

[Ano92g] Anonymous. Index to volume 19, 1992. Science and Public Policy, 19(6):411–412, December 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:Ma

[Ano92h] Anonymous. Meetings. Science and Public Policy, 19(1): CO3, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:Mb

[Ano92i] Anonymous. Meetings. Science and Public Policy, 19(3): co3, June 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:Mc

[Ano92j] Anonymous. Meetings. Science and Public Policy, 19(4): co3, August 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:Md

[Ano92k] Anonymous. Meetings. Science and Public Policy, 19(5): co3, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:PPR

[Ano92l] Anonymous. Programme of Policy Research in Engineering, Science and Technology at Manchester University, UK. Science and Public Policy, 19(3):187– 188, June 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1992:QTA

[Ano92m] Anonymous. Quantifying technological advance: S&T indicators at the OECD — challenges for the 1990s. Science and Public Policy, 19(5):281–290, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1993:D

[Ano93a] Anonymous. Diary. Science and Public Policy, 20(4):CO3, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1993:EELa

[Ano93b] Anonymous. East European literature. Science and Public Policy, 20(1):63–64, February 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1993:EELb

[Ano93c] Anonymous. East European literature. Science and Public Policy, 20(4):293–295, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1993:IS

[Ano93d] Anonymous. Index, SPP 1993.

Science and Public Policy, 20(6): 431–432, December 1993. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1993:Ma

[Ano93e] Anonymous. Meetings. Science and Public Policy, 20(1): CO3, February 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1993:Mb

[Ano93f] Anonymous. Meetings. Science and Public Policy, 20(2): CO3, April 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1993:Mc

[Ano93g] Anonymous. Meetings. Science and Public Policy, 20(3): co3, June 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1993:Md

[Ano93h] Anonymous. Meetings. Science and Public Policy, 20(5): CO3, October 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1993:Me

[Ano93i] Anonymous. Meetings. Science and Public Policy, 20(6): co3, December 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1994:IS

[Ano94a] Anonymous. Index, SPP 1994. Science and Public Policy, 21(6): 427–428, December 1994. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1994:Ma

[Ano94b] Anonymous. Meetings. Science and Public Policy, 21(1): co3, February 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1994:Mb

[Ano94c] Anonymous. Meetings. Science and Public Policy, 21(2): CO3, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1994:Mc

[Ano94d] Anonymous. Meetings. Science and Public Policy, 21(3): cO3–co3, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1994:Md

[Ano94e] Anonymous. Meetings. Science and Public Policy, 21(4): CO3, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1994:Me

[Ano94f] Anonymous. Meetings. Science and Public Policy, 21(5): CO3, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1994:Mf

[Ano94g] Anonymous. Meetings. Science and Public Policy, 21(6): CO3, December 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1995:Da

[Ano95a] Anonymous. Diary. Science and Public Policy, 22(2):co3, April 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1995:Db

[Ano95b] Anonymous. Diary. Science and Public Policy, 22(5):co3, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1995:IS

[Ano95c] Anonymous. Index, SPP 1995.

Science and Public Policy, 22(6): 423–424, December 1995. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1995:Ma

[Ano95d] Anonymous. Meetings. Science and Public Policy, 22(1): CO3, February 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1995:Mb

[Ano95e] Anonymous. Meetings. Science and Public Policy, 22(3): CO3, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1995:Mc

[Ano95f] Anonymous. Meetings. Science and Public Policy, 22(4): co3, August 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1995:Md

[Ano95g] Anonymous. Meetings. Science and Public Policy, 22(6): CO3, December 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1995:NA

[Ano95h] Anonymous. Notes for authors. Science and Public Policy, 22(3): 204, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1999:IS

[Ano99] Anonymous. Index, SPP 1999. Science and Public Policy, 26(6): 451–452, December 1999. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Archibugi:1992:PIT

[Arc92] Daniele Archibugi. Patenting as an indicator of technological innovation: a review. Science and Public Policy, 19(6): 357–368, December 1992. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arthur:1990:SCP

[Art90] Brian Arthur. Six Countries Programme: The international innovation network. Science and Public Policy, 17(5):271– 272, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Atkinson:1993:TLS

[Atk93] Robert D. Atkinson. Ten lessons for successful US federal technology policy-making: learning from the states. Science and Public Policy, 20(1):3–10, February 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Averch:1993:CRE

[Ave93] H. A. Averch. Criteria and rules for evaluating competing R&D megaprojects. Science and Public Policy, 20(2):105–113, April 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Axon:1999:QC

[Axo99] Colin Axon. A question of communication. Science and Public Policy, 26(2):140–141, April 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Baark:1994:ATK

[Baa94] Erik Baark. The arduous transition: key issues of science and technology policy in Mongolia. Science and Public Policy, 21(5): 335–343, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Babu:1999:BPG

[Bab99] D. Senthil Babu. Biotech power games and poor farmers. Science and Public Policy, 26(1): 64–66, February 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bader:1990:CS

[Bad90] Godric Bader. Cartographer of the soul. Science and Public Policy, 17(2):122–123, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Balsiger:1991:ELS

[Bal91] Hans Balsiger. Europe and the large solar system projects. Science and Public Policy, 18(4): 214–218, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Balazs:1992:TST

[Bal92] Katalin Balázs. Transition of the science and technology management system in Hungary. Science and Public Policy, 19(2): 89–97, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Baochen:1991:CGS

[Bao91] Wang Baochen. Chinese government support for science and technology. Science and Public Policy, 18(4):225–233, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barker:1990:SC

[Bar90a] David Barker. Scientific coverup. Science and Public Policy, 17(5):334–335, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barry:1990:CDE

[Bar90b] Andrew Barry. Community and diversity in European technology. Science and Public Policy, 17(6):338–342, December 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barker:1994:SIR

[Bar94a] Katharine Barker. Strengthening the impact of R&D evaluation on policy making: Methodological and organisational considerations. Science and Public Policy, 21(6):405–413, December 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barnett:1994:KTD

[Bar94b] Andrew Barnett. Knowledge transfer and developing countries: The tasks for science and technology in the global perspective 2010. Science and Public Policy, 21(1):2–12, February 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barker:1995:JST

[Bar95] Brendan Barker. Japanese science and technology policy resources on the world wide Web. Science and Public Policy, 22(5):

333–343, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boehmer-Christiansen:1995:RSA

[BC95a] Dr Sonja Boehmer-Christiansen. Reflections on scientific advice and EC transboundary pollution policy. Science and Public Policy, 22(3):195–203, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boehmer-Christiansen:1995:CSB

[BC95b] S. A. Boehmer-Christiansen. Comments on S. Shackley and B. Wynne paper “Global climate change: The mutual construction of an emergent sciencepolicy domain”. Science and Public Policy, 22(6):411–413, December 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Besnosik:1995:MTB

[BdSdOR95] Roberto Iazlovitch Besnosik, Edvaldo Alves de Santana, Adilson de Oliveira, and Maria Tereza Franco Ribeiro. Management of technology in the Brazilian power sector. Science and Public Policy, 22(1): 2–8, February 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bijman:1995:BVC

[BE95] W. Jos Bijman and Christine M. Enzing. Biotechnology and vertical co-ordination in the agrofood chain: a case study of the Dutch potato chain. Science and Public Policy, 22(6): 391–398, December 1995. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bechmann:1993:DFT

[Bec93] Gotthard Bechmann. Democratic function of technology assessment in technology policy decision-making. Science and Public Policy, 20(1):11–16, February 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bell:1999:SS

[Bel99a] David Bell. Secret science. Science and Public Policy, 26(6): 450, December 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bell:1999:T

[Bel99b] David Bell. TechnoBodies. Science and Public Policy, 26(4): 283–284, August 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bell:1999:VP

[Bel99c] David Bell. Virtual politics. Science and Public Policy, 26(1): 66–67, February 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bereano:1999:WIC

[Ber99] Philip L. Bereano. Will informing citizens democratize biotechnology? Science and Public Policy, 26(6):443–444, December 1999. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Bhalla:1990:CCI

[Bha90] A. S. Bhalla. Computerisation in Chinese industry. Science and Public Policy, 17(4):243– 255, August 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Biegelbauer:1995:RHI

[Bie95] Peter Biegelbauer. Rich and highly informative. Science and Public Policy, 22(2):130– 131, April 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Blauwhof:1993:NDT

[BL93] Gertrud Blauwhof and Loet Leydesdorff. New developments in technology studies: evolutionary economics and chaos theory. Science and Public Policy, 20(6): 417–423, December 1993. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Blunden:1990:ACW

[Blu90] Margaret Blunden. Armaments collaboration: what form and what for? Science and Public Policy, 17(3):132–142, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bowonder:1991:ICA

[BM91] B. Bowonder and T. Miyake. Industrial competitiveness: an analysis of the Japanese electronics industry. Science and Public Policy, 18(2):93–110, April 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bowonder:1992:DFL

[BM92] B. Bowonder and Dr T. Miyake.

Development of fuzzy logic technology: an analysis of the Japanese technological innovation process. Science and Public Policy, 19(4):207–220, August 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bondi:1990:LNM

[Bon90] Sir Hermann Bondi. Look, no maths! Science and Public Policy, 17(3):202–203, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bondi:1991:DS

[Bon91] Hermann Bondi. Decades of stimulation. Science and Public Policy, 18(1):7–8, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bowman:1995:FAI

[Bow95] Clem Bowman. Forces affecting industrial research. Science and Public Policy, 22(5): 344–345, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Balazs:1994:AIR

[BP94] Katalin Balázs and Guilherme Ary Plonski. Academicindustry relations in middleincome countries: East Europe and Ibero-America. Science and Public Policy, 21(2):109– 116, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Branscomb:1993:STA

[Bra93] Professor Lewis M. Branscomb. Science and technology advice to the US government: deficiencies and alternatives. Science and Public Policy, 20(2):67–78, April 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bromley:1993:TPG

[Bro93a] D. Allan Bromley. On the theory and practice of giving science advice to government. Science and Public Policy, 20(2): 127–132, April 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brown:1993:MNT

[Bro93b] George E. Brown, Jr. The mother of necessity: technology policy and social equity. Science and Public Policy, 20(6): 411–416, December 1993. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brown:1999:ITR

[Bro99] Nik Brown. Is it time to rethink time? Science and Public Policy, 26(2):141–142, April 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brzoska:1999:EFR

[Brz99] Michael Brzoska. Exploiting foreign research. Science and Public Policy, 26(4):286–287, August 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Buck:1993:PUD

[Buc93] David Buck. Problems with UK defence R&D statistics. Science and Public Policy, 20(1):37–49, February 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Buck:1994:FMG

[Buc94] Dave Buck. Fills a major gap.

Science and Public Policy, 21(3): 196–197, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bud:1990:NCP

[Bud90] Robert Bud. No consensus on priorities. Science and Public Policy, 17(2):124–125, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Budworth:1992:ILC

[Bud92] David Budworth. Innovation literature collected. Science and Public Policy, 19(2):128– 130, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bud:1994:EW

[Bud94] Robert Bud. East of the Wall.

Science and Public Policy, 21(4):

273–274, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Buisseret:1993:RUC

[Bui93] Timothy J. Buisseret. The role of users in collaborative IT research: Experience from the UK’s Information Engineering Advanced Technology Programme (IEATP). Science and Public Policy, 20(5):323– 332, October 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bunting:1999:SEC

[Bun99] Andrea Bunting. Selling energy conservation. Science and Public Policy, 26(5):378–379, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Burgen:1990:F

[Bur90] Sir Arnold Burgen. Foreword. Science and Public Policy, 17(2): 66, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cabral:1991:DAS

[Cab91] Arnoldo Cabral. Developing appropriate S&T policies. Science and Public Policy, 18(2): 138–139, April 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cabrai:1992:AO

[Cab92a] Arnoldo Cabrai. Ambitious objectives. Science and Public Policy, 19(3):189–190, June 1992.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cabrai:1992:HTB

[Cab92b] Arnoldo Cabrai. Hi-tech in Brazil. Science and Public Policy, 19(4):255–256, August 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cabral:1992:STP

[Cab92c] Arnoldo Cabral. Science and technology policy: The Brazilian experience in the aeronautical industry. Science and Public Policy, 19(1):35–42, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cabral:1994:DGV

[Cab94a] Arnoldo Cabral. Developing a global view. Science and Public Policy, 21(5):355–356, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cabral:1994:BWP

[Cab94b]  Regis Cabral.  Biotechnol-

ogy, wheat production, and the Brazilian company for agricultural and livestock research (EMBRAPA), 1970–1990. Science and Public Policy, 21(3): 147–156, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cabrai:1995:SO

[Cab95] Arnoldo Cabrai. Successful and original. Science and Public Policy, 22(4):279–280, August 1995.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Caputo:1990:ISR

[Cap90] Giuseppe Caputo. Individual scholar’s responsibility. Science and Public Policy, 17(2): 73–74, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cardwell:1995:LUA

[Car95] Donald Cardwell. A lesson to us all. Science and Public Policy, 22(2):133–134, April 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cardoza:1999:LIP

[Car99] Guillermo Cardoza. Learning and innovation paths in east Asia. Science and Public Policy, 26(4):259–276, August 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Crow:1991:RLU

[CB91] Michael Crow and Barry Bozeman. R&D laboratories in the USA: Structure, capacity and context. Science and Public Policy, 18(3):165–179, June 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Conceicao:1999:RUK

[CH99] Pedro Conceição and Manuel V. Heitor. On the role of the university in the knowledge economy. Science and Public Policy, 26(1): 37–51, February 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cheese:1994:HU

[Che94] John Cheese. Hunt the user. Science and Public Policy, 21(1): 65–66, February 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Chung:1999:KIP

[Chu99] Sunyang Chung. Korean innovation policies for small and medium-sized enterprises. Science and Public Policy, 26(2): 70–82, April 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Castanos-Lomnitz:1995:UGI

[CL95] Heriberta Castaños-Lomnitz.

University, government and industry in Mexico: the shared dislike of each other. Science and Public Policy, 22(5):325– 332, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Clark:1995:INK

[Cla95]  Norman Clark.  Interactive

nature of knowledge systems: some implications for the Third World. Science and Public Policy, 22(4):249–258, August 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cleland:1995:IUF

[CM95] George Cleland and Donald MacKenzie. The industrial uptake of formal methods in computer science: an analysis and a policy proposal. Science and Public Policy, 22(6):369– 382, December 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Colombo:1992:COS

[Col92]  Umberto Colombo.  Co-

operation in science and technology as a contribution to international security. Science and Public Policy, 19(1):2–6, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Conrad:1993:RST

[Con93] Jobst Conrad. Reflections on science and technology policy advice to government in Germany. Science and Public Policy, 20(2): 97–104, April 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Coopey:1991:UB

[Coo91] Richard Coopey. Uninventing the bomb. Science and Public Policy, 18(4):265–266, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Correa:1994:TTT

[Cor94] Carlos M. Correa. Trends in technology transfer: Implications for developing countries. Science and Public Policy, 21(6): 369–380, December 1994. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cordell:1995:FSR

[Cor95] Arthur J. Cordell. Fuzzy set or readings. Science and Public Policy, 22(5):348–349, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Chubin:1992:ACR

[CR92a] Daryl E. Chubin and Elizabeth M. Robinson. Accounting for the costs of research: Some policy rethinking. Science and Public Policy, 19(3): 181–185, June 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Chubin:1992:HRR

[CR92b] Daryl E. Chubin and Elizabeth M. Robinson. Human resources for the research work force: US indicators and policy choices. Science and Public Policy, 19(6):334–342, December 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Crow:1994:STP

[Cro94] Dr Michael M. Crow. Science and technology policy in the United States: trading in the 1950 model. Science and Public Policy, 21(4):202–212, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dagnino:1993:BLM

[Dag93] Renato Dagnino. To the barracks or into the labs? Military programmes and Brazilian S&T policy. Science and Public Policy, 20(6):389–390, December 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Darier:1999:RDD

[Dar99] Éric Darier. Risks of depolitisation: (un)democratic targets? Science and Public Policy, 26(6): 448–449, December 1999. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

deBruijn:1999:SEC

[dBtH99] J. A. de Bruijn and E. F. ten Heuvelhof. Scientific expertise in complex decision-making processes. Science and Public Policy, 26(3):179–184, June 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

DeMarchi:1995:EPP

[De 95] Bruna De Marchi. Environmental problems, policy decisions and risk communication: What is the role for the social sciences? Science and Public Policy, 22(3): 157–161, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Deckers:1994:TCH

[Dec94] Alain Deckers. Tackling the challenge head on. Science and Public Policy, 21(1):66–67, February 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Deniozos:1994:PRI

[Den94] Dr Dimitris Deniozos. The process of rationalisation in the impact evaluation of RTD restructuring programmes. Science and Public Policy, 21(1): 23–30, February 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ducharme:1992:SAM

[DG92] Louis Marc Ducharme and Fred Gault. Surveys of advanced manufacturing technology. Science and Public Policy, 19(6): 393–399, December 1992. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Djellal:1999:SSR

[DG99] Faridah Djellal and Faïz Gallouj. Services and the search for relevant innovation indicators: a review of national and international surveys. Science and Public Policy, 26(4):218– 232, August 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1991:TWL

[Div91a] Colin Divall. And there was light. Science and Public Policy, 18(6):419–420, December 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1991:IBE

[Div91b] Colin Divall. Inspiration for British engineers? Science and Public Policy, 18(3):200– 203, June 1991. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1992:BFJ

[Div92a]  Colin Divall.  British fail-

ure: Japanese success. Science and Public Policy, 19(5): 329–331, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1992:GFT

[Div92b]  Colin Divall.  Governments

failed training boards. Science and Public Policy, 19(6):407– 408, December 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1992:IDA

[Div92c] Colin Divall. Inadequate defence of assumptions. Science and Public Policy, 19(1):60–61, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1993:HKC

[Div93] Colin Divall. Historical knowledge and contemporary policy. Science and Public Policy, 20(3): 217–218, June 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1994:U

[Div94a] Colin Divall. Not for the uninitiated. Science and Public Policy, 21(2):132, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1994:PNS

[Div94b] Colin Divall. Promise of new social technology. Science and Public Policy, 21(1):67–68, February 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1994:RP

[Div94c] Colin Divall. Reflections on the past. Science and Public Policy, 21(4):274–276, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1995:ARP

[Div95a] Colin Divall. Asymmetric relations of power. Science and Public Policy, 22(1):66–67, February 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1995:GTP

[Div95b]  Colin Divall.  Guides for

technology-policy commentators. Science and Public Policy, 22(2):132–133, April 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

deJong:1999:ICD

[dJ99] Martin de Jong. Institutionalised criticism: the demonopolisation of scientific advising. Science and Public Policy, 26(3): 193–199, June 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Durrenberger:1999:IAF

[DKB99] Gregor Dürrenberger, Hans Kastenholz, and Jeannette Behringer. Integrated assessment focus groups: Bridging the gap between science and policy? Science and Public Policy, 26(5): 341–349, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

delaMothe:1992:PNS

[dlM92] John de la Mothe. The political nature of science and indicators. Science and Public Policy, 19(6):401–406, December 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

delaMothe:1994:SY

[dlM94] John de la Mothe. The snows of yesteryear. Science and Public Policy, 21(6):421–423, December 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

delaMothe:1995:WWD

[dlM95] John de la Mothe. Which way did they go? Science and Public Policy, 22(6):421–422, December 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

delaMothe:1994:CEN

[dlMP94] John de la Mothe and Gilles Paquet. Circumstantial evidence: a note on science policy in Canada. Science and Public Policy, 21(4):261–268, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Doms:1992:THT

[DM92] Mark E. Doms and Robert H.

McGuckin. Trade in high technology products. Science and Public Policy, 19(6):343– 346, December 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

DeMarchi:1993:SRD

[DN93] Mario De Marchi and Giovanni Napolitano. Some revised definitions of applied research and experimental development. Science and Public Policy, 20(4): 281–284, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

deOliveira:1991:ETP

[dO91] Adilson de Oliveira. Energy technology policy in developing countries. Science and Public Policy, 18(3):156–164, June 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Doi:1994:FSR

[Doi94] Noriyuki Doi. Firm size and R&D activity in Japanese manufacturing industries. Science and Public Policy, 21(5):295– 308, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dosi:1991:PET

[Dos91] Giovanni Dosi. Perspectives on evolutionary theory. Science and Public Policy, 18(6):353– 361, December 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Drenth:1990:SUT

[Dre90] P. J. D. Drenth. “inter utrumque tene”: the two responsibilities of the scientist. Science and Public Policy, 17(2):75–77, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Damiani:1995:MI

[dSD95] José Henrique de Sousa Damiani. Marketing IT. Science and Public Policy, 22(4):274– 275, August 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ducatel:1990:VCV

[Duc90] Dr Ken Ducatel. Vans are not commercial vehicles. Science and Public Policy, 17(2): 123–124, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dufour:1993:BMD

[Duf93a] Paul Dufour. Balance of merits and dangers of S&T. Science and Public Policy, 20(6): 424–425, December 1993. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dufour:1993:MH

[Duf93b] Paul Dufour. A man with hope. Science and Public Policy, 20(4): 286–288, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dufour:1995:LCC

[Duf95a] Paul Dufour. Lucky country, cold colony. Science and Public Policy, 22(5):351–352, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dufour:1995:SP

[Duf95b]  Paul Dufour.  A sense of

progress. Science and Public Policy, 22(4):275–276, August 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dufour:1995:TCW

[Duf95c] Paul Dufour. Technology and competitiveness for what? Science and Public Policy, 22(6): 419–421, December 1995. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dufour:1995:TA

[Duf95d] Paul Dufour. A tenure of accomplishment. Science and Public Policy, 22(3):211–212, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Durant:1999:PTA

[Dur99] John Durant. Participatory technology assessment and the democratic model of the public understanding of science. Science and Public Policy, 26(5): 313–319, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Edgerton:1991:IT

[Edg91] David Edgerton. Invention of tradition. Science and Public Policy, 18(4):267–268, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Edmonds:1994:PIA

[Edm94] Dr Martin Edmonds. Potentially important but acutely disappointing. Science and Public Policy, 21(6):417–418, December 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Edwards:1999:SEP

[Edw99] Arthur Edwards. Scientific expertise and policy-making: the intermediary role of the public sphere. Science and Public Policy, 26(3):163–170, June 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Elzen:1990:WIN

[EES90] Boelie Elzen, Bert Enserink, and Wim A. Smit. Weapon innovation: networks and guiding principles. Science and Public Policy, 17(3):171–193, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Efana:1993:IAS

[Efa93] Lawrence U. B. Efana. Internationalisation of academic and scientific contacts: Aspects of Finnish scholarly exchanges. Science and Public Policy, 20(5): 351–359, October 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Etzkowitz:1999:PET

[EG99] Henry Etzkowitz and Magnus Gulbrandsen. Public entrepreneur: the trajectory of United States science, technology and industrial policy. Science and Public Policy, 26(1): 53–62, February 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ekins:1992:EHG

[Eki92] Paul Ekins. Environmental head of government. Science and Public Policy, 19(1):61–63, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Etzkowitz:1992:AUB

[EKNU92] Henry Etzkowitz, Carol Kemelgor, Michael Neuschatz, and Brian Uzzi. Athena unbound: Barriers to women in academic science and engineering. Science and Public Policy, 19(3): 157–179, June 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Etzkowitz:1999:WTH

[EL99] Henry Etzkowitz and Loet Leydesdorff. Whose Triple Helix? Science and Public Policy, 26(2): 138–139, April 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Etzkowitz:1998:WPD

[EMT98] Henry Etzkowitz, José Manoel C. Mello, and Branca Regina Cantisano Terra. When path dependencies collide: the evolution of innovation policy in the State of Rio de Janeiro, Brazil. Science and Public Policy, 25(6):365–371, December 1998. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Endersby:1999:TDB

[End99a] Jim Endersby. Three different books trying to get out. Science and Public Policy, 26(6): 447–448, December 1999. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Endersby:1999:WS

[End99b] Jim Endersby. What science is not. … Science and Public Policy, 26(3):211–212, June 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Enzing:1990:GRI

[ES90] Christien M. Enzing and Ruud E. H. M. Smits. Government’s role in the internationalisation of R&D: The increasing demand for strategic decision-making in science and technology. Science and Public Policy, 17(5): 287–292, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Etzkowitz:1994:TCI

[Etz94] Henry Etzkowitz. Technology centers and industrial policy: the emergence of the interventionist state in the USA. Science and Public Policy, 21(2): 79–87, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Edmonds:1990:UUD

[EUH90] Martin Edmonds, Matthew Uttley, and George Hayhurst. UK and US dependence on foreign technology in defence research and development. Science and Public Policy, 17(3): 157–170, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Evans:1999:MPC

[Eva99] Dr Robert Evans. Model policies for climate change and trasport. Science and Public Policy, 26(6): 444–445, December 1999. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Everest:1992:ED

[Eve92] D. A. Everest. Environmental diplomacy. Science and Public Policy, 19(3):193–194, June 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Faulkner:1995:GBI

[Fau95] Wendy Faulkner. Getting behind industry-public sector research linkage: a novel research design. Science and Public Policy, 22(5):282–294, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fernandez:1995:DDE

[Fer95] Dr Susana Aguilar Fernández. Differences and dynamics in European Union environmental policy. Science and Public Policy, 22(3):189–194, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Faucher:1993:PDM

[FF93] Philippe Faucher and Kevin Fitzgibbons. Public demand and the management of technological risk in large-scale projects. Science and Public Policy, 20(3): 173–185, June 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Finney:1999:EPC

[Fin99] Colin Finney. Extending public consultation via the Internet: The experience of the UK Advisory Committee on Genetic Testing electronic consultation. Science and Public Policy, 26(5): 361–373, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fischer:1999:TDD

[Fis99] Frank Fischer. Technological deliberation in a democratic society: The case for participatory inquiry. Science and Public Policy, 26(5):294–302, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fitzgerald:1993:MCA

[Fit93] Lawrence Fitzgerald. Museums as a cultural attraction. Science and Public Policy, 20(4): 288–289, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ford:1991:EES

[FL91] Glyn Ford and Gordon Lake.

Evolution of European science and technology policy. Science and Public Policy, 18(1):38–50, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forje:1990:RCH

[For90] John W. Forje. Reawakening cultural heritage. Science and Public Policy, 17(1):58–64, February 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forje:1995:KEM

[For95a] Dr John W. Forje. Knowledge and ethics in the modern world. Science and Public Policy, 22(6): 417–418, December 1995. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forje:1995:DTP

[For95b] John W. Forje. Development and technology policy in a changing society: challenges for the African political structures. Science and Public Policy, 22(2): 121–129, April 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Freeman:1991:TPQ

[Fre91] Chris Freeman. Technology, progress and the quality of life.

Science and Public Policy, 18(6): 407–418, December 1991. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fryers:1990:UF

[Fry90] G. R. Fryers. Utopian fallacy. Science and Public Policy, 17(6): 387–388, December 1990. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Galhardi:1994:BPB

[Gal94] Regina M. A. A. Galhardi. Brazilian policy for biotechnology: a critical review. Science and Public Policy, 21(6): 395–403, December 1994. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Garfield:1991:CIS

[Gar91] Dr Eugene Garfield. Contribution of the international science policy foundation. Science and Public Policy, 18(1):6, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gaillard:1993:FAA

[GB93] Jacques Gaillard and Lawrence Busch. French and American agricultural science for the Third World. Science and Public Policy, 20(4):222–234, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ganz-Brown:1999:PPF

[GB99] Dr Carole Ganz-Brown. Patent policies to fine tune commercialization of government-sponsored university research. Science and Public Policy, 26(6):403– 414, December 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Geffen:1995:RIE

[Gef95] Charlette A. Geffen. Radical innovation in environmental technologies: the influence of federal policy. Science and Public Policy, 22(5):313–323, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

George:1993:CAP

[Geo93a] A. D. George. Coherent aviation policy required. Science and Public Policy, 20(3):215– 217, June 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

George:1993:DS

[Geo93b] A. D. George. Definitive study.

Science and Public Policy, 20(6): 429–430, December 1993. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

George:1994:WFI

[Geo94] A. D. George. Wealth of fascinating information. Science and Public Policy, 21(2):131– 132, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gibbs:1993:EE

[Gib93] David Gibbs. Environmental economics. Science and Public Policy, 20(6):428–429, December 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gibbs:1994:VFI

[Gib94] David Gibbs. Vision of the future ignores problems. Science and Public Policy, 21(6):423– 424, December 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goodwin:1999:PAC

[GJ99] Dr Miriam Goodwin and Ron Johnston. The place of absorptive capacity in national innovation systems: the case of Australia. Science and Public Policy, 26(2):83–90, April 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gagne:1992:QSI

[GL92] Jean Gagné and Michel Leclerc. Quebec’s science indicators: a survey. Science and Public Policy, 19(5):303–310, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Glover:1995:DBM

[Glo95] Ian Glover. Deserves to become a minor classic. Science and Public Policy, 22(2):135– 136, April 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1990:SCE

[Gol90] Maurice Goldsmith. ”one culture” embraces “two”. Science and Public Policy, 17(1):2, February 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1991:ROI

[Gol91] Maurice Goldsmith. Reshaping our ideological map. Science and Public Policy, 18(4):206, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1992:RRS

[Gol92] David A. Goldsmith. The rise and rise of SKY. Science and Public Policy, 19(3):190– 193, June 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goncalves:1995:SEE

[Gon95] Maria Eduarda Gonçalves. Scientific expertise and European Community regulatory processes. Science and Public Policy, 22(3):183–187, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gooday:1993:RUS

[Goo93] Graeme Gooday. Reconstructing US science teaching. Science and Public Policy, 20(5): 366–367, October 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gooday:1994:EBS

[Goo94a] Graeme Gooday. Evolutionary biology to social policy. Science and Public Policy, 21(5): 352–354, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gooday:1994:RUR

[Goo94b] Graeme Gooday. Rough, underresearched draft. Science and Public Policy, 21(3):193–194, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gordon:1992:TC

[Gor92] John Gordon. Time for change. Science and Public Policy, 19(2): 126–127, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Govindarajulu:1990:EEI

[Gov90] V. Govindarajulu. Evaluation of effectiveness of India’s public policy options. Science and Public Policy, 17(6):349–362, December 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Glasner:1999:DFB

[GR99] Peter Glasner and Harry Rothman. Does familiarity breed concern? Bench scientists and the human genome mapping project. Science and Public Policy, 26(4):233–240, August 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Green:1992:EIA

[Gre92] Dr Ken Green. Economic issues avoided. Science and Public Policy, 19(1):59–60, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Greenstein:1993:GBC

[Gre93] Shane Greenstein. Going by the book: the costs and benefits of procedural rules in federal computer procurement. Science and Public Policy, 20(3): 157–171, June 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Groenewegen:1990:RD

[Gro90a] Dr Peter Groenewegen. Richness in diversity. Science and Public Policy, 17(1):56–58, February 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Groenewegen:1990:II

[Gro90b] Peter Groenewegen. International interconnections. Science and Public Policy, 17(6):388– 390, December 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Grobicki:1994:FDS

[Gro94] Dr Ania Grobicki. The formulation of a democratic science and technology policy in South Africa: the ANC policy process 1990–1992. Science and Public Policy, 21(4):213–220, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1991:EST

[Gum91] Philip Gummett. The evolution of science and technology policy: a UK perspective. Science and Public Policy, 18(1):31–37, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1992:DT

[Gum92] Philip Gummett. Destabilising technologies. Science and Public Policy, 19(6):409–410, December 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1993:IG

[Gum93] Dr Philip Gummett. Intelligence gathering. Science and Public Policy, 20(2):134–135, April 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1994:LTE

[Gum94] Philip Gummett. Lacking a ‘theory of expert advice’. Science and Public Policy, 21(1):64–65, February 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1995:HRP

[Gum95a] Philip Gummett. Highly readable, popular account. Science and Public Policy, 22(2):131– 132, April 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1995:MU

[Gum95b] Philip Gummett. Myth uncovered. Science and Public Policy, 22(1):65–66, February 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Garfield:1992:CDT

[GWD92] Eugene Garfield and Alfred Welljams-Dorof. Citation data: their use as quantitative indicators for science and technology evaluation and policy-making. Science and Public Policy, 19(5): 321–327, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hane:1995:CFA

[Han95] Gerald J. Hane. Clearing the fog around R&D consortia in Japan. Science and Public Policy, 22 (2):85–94, April 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Harbor:1990:ACM

[Har90] Bernard Harbor. Arms conversion and military-civilian technological synergy. Science and Public Policy, 17(3):194–200, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1990:PG

[Hay90] George Hay. Plugging the gap.

Science and Public Policy, 17(5): 335–336, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1992:GB

[Hay92] George Hay. Government beware! Science and Public Policy, 19(1):63–64, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hayward:1993:ICS

[Hay93] Keith Hayward. International collaboration in space: The case of the International Space Station, Freedom. Science and Public Policy, 20(5):333–341, October 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Healey:1990:ISI

[Hea90] Peter Healey. Introduction to special issue on defence R&D. Science and Public Policy, 17(3): 130–131, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Healey:1994:EAI

[Hea94]  Peter Healey.  Evaluating

academic-industry relations: the establishment and contribution of an international study group. Science and Public Policy, 21 (2):70–71, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hempstead:1994:TFT

[Hem94] Dr Colin A. Hempstead. Too fragmented, too eulogistic. Science and Public Policy, 21(6): 419–420, December 1994. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hempstead:1995:EHT

[Hem95] Dr Colin A. Hempstead. Elucidating history of technology. Science and Public Policy, 22(1): 71–72, February 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hennen:1999:PTA

[Hen99] Leonhard Hennen. Participatory technology assessment: a response to technical modernity? Science and Public Policy, 26(5): 303–312, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Herman:1994:COR

[Her94] Ros Herman. Co-ordinating research and technological development in Europe: Lessons from the German system. Science and Public Policy, 21(6):358– 368, December 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hinder:1990:PSR

[HH90] Susan Hinder and Peter Halfpenny. Potential for sharing research equipment. Science and Public Policy, 17(2):105– 110, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Horlick-Jones:1995:GE

[HJ95] Tom Horlick-Jones. Guest editorial. Science and Public Policy, 22(3):138, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Horlick-Jones:1995:CSE

[HJD95] Tom Horlick-Jones and Bruna De Marchi. The crisis of scientific expertise in fin de siècle Europe. Science and Public Policy, 22(3):139–145, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hartley:1990:ICA

[HM90] Keith Hartley and Dr Stephen Martin. International collaboration in aerospace. Science and Public Policy, 17(3):143– 151, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holbrook:1991:ISE

[Hol91] J. A. D. Holbrook. The influence of scale effects on international comparisons of R&D expenditures. Science and Public Policy, 18(4):259–262, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holbrook:1992:BIS

[Hol92a] J. A. D. Holbrook. Basic indicators of scientific and technological performance. Science and Public Policy, 19(5):267– 273, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holbrook:1992:WMS

[Hol92b] J. A. D. Holbrook. Why measure science? Science and Public Policy, 19(5):262–266, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holbrook:1995:FCC

[Hol95a] Adam Holbrook. A five-cent cigar for the policy-maker. Science and Public Policy, 22(5): 345–347, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holbrook:1995:GRV

[Hol95b] Adam Holbrook. A good review from Van Dieman’s Land. Science and Public Policy, 22(4): 277–279, August 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holbrook:1995:OOG

[Hol95c] Adam Holbrook. OECD not the only game in town. Science and Public Policy, 22(3): 213–214, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Honig:1991:SPS

[Hon91] Carlye Honig. Science policy support group, London. Science and Public Policy, 18(2): 135–136, April 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Honig:1993:GD

[Hon93] Carlye Honig. Gender discrimination. Science and Public Policy, 20(1):60–61, February 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hooper:1993:SRS

[Hoo93] Paul D. Hooper. Sustainability requires ‘something extra’. Science and Public Policy, 20(1): 61–62, February 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hoppe:1999:PAS

[Hop99] Robert Hoppe. Policy analysis, science and politics: from ‘speaking truth to power’ to ‘making sense together’. Science and Public Policy, 26(3): 201–210, June 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Horrocks:1994:SSM

[Hor94] Sally M. Horrocks. Surplus skimmed milk to international giant. Science and Public Policy, 21(6):420–421, December 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Horning:1999:CPF

[Hör99] Georg Hörning. Citizens’ panels as a form of deliberative technology assessment. Science and Public Policy, 26(5):351– 359, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Howells:1990:GRD

[How90] J. Howells. The globalisation of research and development: a new era of change? Science and Public Policy, 17(5):273– 285, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hartley:1990:DRC

[HS90] Keith Hartley and Dr John Singleton. Defence R&D and crowding-out. Science and Public Policy, 17(3):152–156, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hohn:1991:PDC

[HS91] Hans-Willy Hohn and Volker Schneider. Path-dependency and critical mass in the development of research and technology: a focused comparison. Science and Public Policy, 18(2): 111–122, April 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hudson:1999:FS

[Hud99] Robert G. Hudson. Fallibility of science. Science and Public Policy, 26(4):284–286, August 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ingelstam:1994:IEP

[Ing94] Lars Ingelstam. Issues of expansion: Perspectives on IT policy in Sweden. Science and Public Policy, 21(6):387–394, December 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irvine:1991:PIS

[Irv91] John Irvine. Promoting innovation in scientific instruments: Some lessons for government policy. Science and Public Policy, 18(3):181–195, June 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1991:CCS

[Irw91a] Alan Irwin. Call for critical science journalism. Science and Public Policy, 18(2):137– 138, April 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1991:TR

[Irw91b] Alan Irwin. For the troubled regulator. Science and Public Policy, 18(3):199–200, June 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1991:PP

[Irw91c] Alan Irwin. Playfulness and provocation. Science and Public Policy, 18(6):420–422, December 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1992:LCO

[Irw92a] Alan Irwin. Lack of co-ordinated policy. Science and Public Policy, 19(2):127–128, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1992:WPG

[Irw92b] Alan Irwin. Where are the pressure groups? Science and Public Policy, 19(4):256–257, August 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1993:PNR

[Irw93] Alan Irwin. Pollution no respecter of boundaries. Science and Public Policy, 20(5):362– 363, October 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ionescu-Sisesti:1994:RRS

[IS94] Ileana Ionescu-Sisesti. Restructuring the R&D system in Romania. Science and Public Policy, 21(1):55–60, February 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Isnor:1993:FBP

[Isn93] Richard Isnor. Federal biotechnology policy in Canada — relative success or ineffectiveness? Science and Public Policy, 20(1): 17–25, February 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jaffe:1992:ABR

[Jaf92] Walter R. Jaffé. Agricultural biotechnology research and development investment in some Latin American countries. Science and Public Policy, 19(4): 229–240, August 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jasinski:1991:DPI

[Jas91a] Dr Andrzej H. Jasinski. Dilemmas of public innovation policy in Poland in the transition to a market economy. Science and Public Policy, 18(2):131– 134, April 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jasinski:1991:RCP

[Jas91b] Dr Andrzej H. Jasinski. Recent changes in the Polish R&D system. Science and Public Policy, 18(4):245–249, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jasinski:1994:VST

[Jas94] Dr Andrzej H. Jasinski. Viewpoint: Science and technology policy and changes in Polish industry in the transition period. Science and Public Policy, 21(3): 188–192, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Joss:1999:CCP

[JB99] Simon Joss and Arthur Brownlea. Considering the concept of procedural justice for public policy- and decision-making in science and technology. Science and Public Policy, 26(5): 321–330, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jennings:1999:CNP

[Jen99a] Richard C. Jennings. Changes needed to preserve biodiversity. Science and Public Policy, 26(5): 379–380, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jennings:1999:ICE

[Jen99b] Richard C. Jennings. Interesting collection of essays. Science and Public Policy, 26(2): 142–144, April 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jevons:1992:EMT

[Jev92a] Fred Jevons. Exposing muddled thinking. Science and Public Policy, 19(6):408–409, December 1992. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Jevons:1992:TMA

[Jev92b] Fred Jevons. Truth must not be abandoned. Science and Public Policy, 19(4):254–255, August 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jevons:1993:CLA

[Jev93] Fred Jevons. The co-location assumption and models of innovation. Science and Public Policy, 20(1):51–56, February 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Johnston:1993:DES

[Joh93] Professor Ron Johnston. Designing an effective science and technology policy advisory system for the Australian government. Science and Public Policy, 20 (2):87–96, April 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Johnston:1999:BDT

[Joh99a] Sean F. Johnston. A big dish tastefully presented. Science and Public Policy, 26(3):215– 216, June 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Johnston:1999:TT

[Joh99b] Sean F. Johnston. Two tribes.

Science and Public Policy, 26(5): 375–376, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jones:1993:LEE

[Jon93] Oswald Jones. Lack of empirical evidence. Science and Public Policy, 20(5):364–366, October 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Joseph:1994:INA

[Jos94] Richard Joseph. Important and neglected area of analysis. Science and Public Policy, 21(4): 278–279, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Joss:1999:PPS

[Jos99] Simon Joss. Public participation in science and technology policy- and decision-making — ephemeral phenomenon or lasting change? Science and Public Policy, 26(5):290–293, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kaldor:1991:EEA

[Kal91] Mary Kaldor. Eastern Europe: aspirations and reality. Science and Public Policy, 18(6):369– 373, December 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Karel:1994:EMA

[Kar94]  Müller Karel.  Extending

Meske’s analysis of East German R&D to other East European countries. Science and Public Policy, 21(1):61–63, February 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Katz:1991:IRP

[Kat91] Jorge M. Katz. Industrial restructuring, public expenditure and social equity: the experience of Argentina. Science and Public Policy, 18(6):375–378, December 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Keith:1992:TII

[Kei92] Smith Keith. Technological innovation indicators: experience and prospects. Science and Public Policy, 19(6):383–392, December 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kellermann:1992:FD

[Kel92] E. W. Kellermann. Fascinating and dangerous. Science and Public Policy, 19(5):331– 332, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kemp:1993:IST

[Kem93] Stewart Kemp. Is it safe to transport radioactive materials? Science and Public Policy, 20(1): 57–58, February 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Khan:1991:GWI

[Kha91] Dr R. R. Khan. Global warming — implications and proposals. Science and Public Policy, 18(6):424–426, December 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Klochko:1993:IMV

[KI93] Yuri Klochko and Nina Issakova. Intellectual migration: a view from Ukraine. Science and Public Policy, 20(6):405–409, December 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kienitz:1995:ANN

[Kie95] Karl Heinz Kienitz. Application of neural networks. Science and Public Policy, 22(3): 214–215, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

King:1991:GTR

[Kin91] Alexander King. The great transition restated. Science and Public Policy, 18(1):15–22, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kinder:1999:SAR

[KKM99] Tony Kinder, Matthias Klaes, and Alfonso Molina. Sociotechnical alignment in the rise and evolution of a telemedicine constituency in Scotland. Science and Public Policy, 26(6):415– 435, December 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kneen:1990:BP

[Kne90] Dr Peter Kneen. Before perestroika. Science and Public Policy, 17(4):264–266, August 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kneen:1993:SSL

[Kne93] Dr Peter Kneen. The Soviet scientific legacy: some differences of interpretation. Science and Public Policy, 20(4): 251–260, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kneller:1999:IPR

[Kne99] Robert Kneller. Intellectual property rights and universityindustry technology transfer in Japan. Science and Public Policy, 26(2):113–124, April 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kodama:1991:CGP

[Kod91]  Fumio Kodama.  Changing

global perspective: Japan, the USA and the new industrial order. Science and Public Policy, 18(6):385–392, December 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Koizumi:1993:HTP

[Koi93] Kenkichiro Koizumi. Historical turning points in Japanese joint research policy. Science and Public Policy, 20(5):313– 322, October 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Koj:1990:CRS

[Koj90] Aleksander Koj. Complex responsibilities of scientists in the contemporary world. Science and Public Policy, 17(2):78– 81, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kostoff:1994:ARI

[Kos94] Dr Ronald N. Kostoff. Assessing research impact: US government retrospective and quantitative approaches. Science and Public Policy, 21(1):13–22, February 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kash:1993:TST

[KR93] Don E. Kash and Robert W. Rycroft. Two streams of technological innovation: implications for policy. Science and Public Policy, 20(1):27–36, February 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Krige:1992:PSP

[Kri92] John Krige. Political and social problems revealed. Science and Public Policy, 19(4): 259–260, August 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kim:1994:IPP

[KRY94] Sang-Gon Kim, Kong-Kyun Ro, and Pyung-Il Yu. Intellectual property protection policy and technology capability. Science and Public Policy, 21(2):121– 130, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kristapsons:1995:CLR

[KT95] Dr Janis Kristapsons and Dr Erika Tjunina. Changes in the Latvian research system. Science and Public Policy, 22(5): 305–312, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Krohn:1994:SLS

[KW94] Wolfgang Krohn and Johannes Weyer. Society as a laboratory: the social risks of experimental research. Science and Public Policy, 21(3):173–183, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lacy:1995:SEC

[Lac95]  William B. Lacy.  Socio-

economic context and policy strategies for US public agricultural sciences. Science and Public Policy, 22(4):239–247, August 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lake:1995:UST

[Lak95] Gordon Lake. Utilisation of scientific and technical expertise in a European policy context. Science and Public Policy, 22(3): 162–168, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lee:1991:SED

[Lee91] Chong-Ouk Lee. Stages of economic development and technology policy: the experience of Korea. Science and Public Policy, 18(4):219–224, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Levidow:1999:EU

[Lev99] Les Levidow. Expert understandings. Science and Public Policy, 26(4):287–288, August 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lichfield:1991:MGI

[Lic91] Nathaniel Lichfield. Maurice Goldsmith: an introduction: Nathaniel Lichfield opens this set of essays in honour of Dr Maurice Goldsmith. Science and Public Policy, 18(1):2–4, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lin:1990:NNS

[Lin90] Li Guang Lin. The National Natural Science Foundation of China: Its background, present, and guiding principles. Science and Public Policy, 17(4):235– 241, August 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lisle:1995:SDB

[Lis95] Edmond Lisle. Scientific discovery — for business, for society — at any price? Science and Public Policy, 22(4): 267–273, August 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Liyanage:1993:CPS

[Liy93] Shantha Liyanage. Changing perspectives of science and technology development in developing countries. Science and Public Policy, 20(4):235–244, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Liyanage:1994:STP

[Liy94] Shantha Liyanage. Science and technology policy Asian network (STEPAN). Science and Public Policy, 21(5):344–346, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lipsett:1995:BYN

[LL95] Morley S. Lipsett and Richard G.

Lipsey. Benchmarks, yardsticks and new places to look for industrial innovation and growth. Science and Public Policy, 22(4): 259–265, August 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Link:1999:WSE

[LL99] Albert N. Link and Jamie R.

Link. Women in science: an exploratory analysis of trends in the United States. Science and Public Policy, 26(6):437– 442, December 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lloyd:1991:SPD

[Llo91] Sir Ian Lloyd. Science, parliament and democracy. Science and Public Policy, 18(1):57–63, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Leclerc:1992:SCO

[LOFM92] Michel Leclerc, Yoshiko Okubo, Luiz Frigoletto, and Jean-François Miquel. Scientific cooperation between Canada and the European Community. Science and Public Policy, 19(1): 15–24, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Longhurst:1993:PIC

[Lon93] Dr Jim Longhurst. Playing it cool. Science and Public Policy, 20(4):285–286, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Longhurst:1994:SEE

[Lon94]  Dr James Longhurst.  Sta-

ble economy essential for environment. Science and Public Policy, 21(5):354–355, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Levidow:1991:GBG

[LT91] Les Levidow and Joyce Tait. The greening of biotechnology: GMOs as environment-friendly products. Science and Public Policy, 18(5):271–280, October 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Levidow:1993:ABR

[LT93] Les Levidow and Joyce Tait.

Advice on biotechnology regulation: the remit and composition of Britain’s ACRE. Science and Public Policy, 20(3): 193–209, June 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lust:1990:FFY

[Lüs90] Reimar Lüst. Forty four years after the atomic bomb: the physicist’s responsibility. Science and Public Policy, 17(2):82–84, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lust:1991:MS

[Lüs91] Reimar Lüst. Man in space. Science and Public Policy, 18(4): 207–213, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mackay:1991:SC

[Mac91] Alan Mackay. The science critic.

Science and Public Policy, 18(1): 11–14, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

MacKenzie:1994:CRA

[Mac94] Donald MacKenzie. Computerrelated accidental death: an empirical exploration. Science and Public Policy, 21(4):233– 248, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mahroum:1999:CHS

[Mah99] Sami Mahroum. Competing for the highly skilled: Europe in perspective. Science and Public Policy, 26(1):17–25, February 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Maricic:1990:SSE

[Mar90] Siniša Maricic. Science-and-society in Europe before World War II. Science and Public Policy, 17(4):257–260, August 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Martin:1992:SNV

[Mar92]  Brian Martin.  Science for

non-violent struggle. Science and Public Policy, 19(1):55–58, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marr:1993:LTI

[Mar93] Liz Marr. Linking transport and IT. Science and Public Policy, 20(2):135–136, April 1993. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marques:1994:BUC

[Mar94a] Dr Marilia Bernardes Marques. Brazil–US controversy on the impact of patenting in biotechnology: some relevant questions for pharmaceuticals. Science and Public Policy, 21(3):165– 172, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marsden:1994:BTC

[Mar94b] Dr Ben Marsden. From biblical times to 20th century. Science and Public Policy, 21(3): 198–200, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marsden:1995:SSW

[Mar95] Dr Ben Marsden. Seamed and seamless webs. Science and Public Policy, 22(1):67–69, February 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mayor:1991:IP

[May91] Federico Mayor. Internationalist and poet. Science and Public Policy, 18(1):5, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mayda:1999:PRT

[May99] Dr Jaro Mayda. Policy R&D:

Toward a better bridge between knowledge and decision making. Science and Public Policy, 26(6): 395–402, December 1999. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mentzel:1999:TTP

[Men99] Maarten Mentzel. Think tanks, policy-making, and a Dutch advisory council. Science and Public Policy, 26(3):171–178, June 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meske:1993:REG

[Mes93] Werner Meske. The restructuring of the East German research system — a provisional appraisal. Science and Public Policy, 20(5):298–312, October 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Metcalfe:1993:APM

[Met93] J. S. Metcalfe. Agenda for policy making. Science and Public Policy, 20(3):218–220, June 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meyer:1990:LBN

[Mey90] Philippe Meyer. Liberty: a biological necessity. Science and Public Policy, 17(2):85–87, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Maital:1994:RBS

[MFGK94] Shlomo Maital, Amnon Frenkel, Hariolf Grupp, and Knut Koschatzky. Relation between scientific and technological excellence and export performance: theoretical model and empirical test for EC countries. Science and Public Policy, 21(3): 138–146, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mardapitta-Hadjipandeli:1992:ITS

[MHN92] Lida Mardapitta-Hadjipandeli and Eri Nicolaides. Information technology in a small country: Potential and practice in Cyprus. Science and Public Policy, 19(1):25–34, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Martin:1992:TGS

[MI92] Ben R. Martin and John Irvine.

Trends in government spending on academic and related research: an international comparison. Science and Public Policy, 19(5):311–319, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Martin:1990:TUG

[MII90] Ben R. Martin, John Irvine, and Phoebe A. Isard. Trends in UK government spending on academic and related research: a comparison with F R Germany, France, Japan, The Netherlands and USA. Science and Public Policy, 17(1):3–13, February 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mills:1992:MTH

[Mil92] Richard Mills. More than a history. Science and Public Policy, 19(5):328–329, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Martin:1990:RTO

[MIN^{+}90] Ben R. Martin, John Irvine, Francis Narin, Kimberly A. Stevens, and Chris Sterritt. Recent trends in the output and impact of British science. Science and Public Policy, 17(1): 14–26, February 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Molina:1990:ENC

[Mol90] Alfonso H. Molina. Emerging neural computing in the USA, Japan and UK/Europe. Science and Public Policy, 17(6): 363–371, December 1990. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Molina:1992:EIS

[Mol92] Alfonso H. Molina. Europe’s IT strategy moves closer to the users: The Open Microprocessor Systems Initiative (OMI). Science and Public Policy, 19(3): 145–156, June 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Molina:1994:UEL

[Mol94] Alfonso H. Molina. Understanding the emergence of a largescale European initiative in technology. Science and Public Policy, 21(1):31–41, February 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mort:1994:OS

[Mor94] Maggie Mort. Overburdened with statistics. Science and Public Policy, 21(4):276–277, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Moran:1995:LC

[Mor95] Michael Moran. Lack of cohesion. Science and Public Policy, 22(3):210–211, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mott:1991:SP

[Mot91] Sir Nevill Mott. Saving the planet. Science and Public Policy, 18(5):328–329, October 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meyer-Stamer:1992:EBI

[MS92] Jörg Meyer-Stamer. The end of Brazil’s informatics policy. Science and Public Policy, 19(2): 99–110, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meyer-Stamer:1995:NDT

[MS95] Jörg Meyer-Stamer. New departures for technology policy in Brazil. Science and Public Policy, 22(5):295–304, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mansell:1999:OKB

[MSdM99] Dr Robin Mansell, Dr W. Edward Steinmueller, and Uta Wehn de Montalvo. Opportunities for knowledge-based development: capabilities, infrastructure, investment and policy. Science and Public Policy, 26(2):91– 100, April 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Munoz:1990:CSP

[Muñ90] Emilio Muñoz. Common scientific policy in Europe: The COST experience. Science and Public Policy, 17(6):343– 348, December 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Munson:1995:RAL

[Mun95] Abby Munson. Risk associated with and liability arising from, releases of genetically manipulated organisms into the environment. Science and Public Policy, 22(1):51–63, February 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mundell:1999:FCF

[Mun99] Ian Mundell. Food comes first. Science and Public Policy, 26(5): 374–375, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mytelka:1990:NMI

[Myt90] Dr Lynn Krieger Mytelka. New modes of international competition: The case of strategic partnering in R&D. Science and Public Policy, 17(5):296– 302, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Narin:1992:TIS

[NAS92] Francis Narin, Michael B. Albert, and Vincent M. Smith. Technology indicators in strategic planning. Science and Public Policy, 19(6):369–381, December 1992. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Nascimento:1995:RUS

[Nas95] Dr Cairo L. Nascimento, Jr. Rethinking utility services. Science and Public Policy, 22(3): 215–216, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nelson:1991:RFD

[Nel91] Richard R. Nelson. The role of firm differences in an evolutionary theory of technical advance. Science and Public Policy, 18(6): 347–352, December 1991. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nelis:1999:RCI

[Nel99a] Annemiek Nelis. Richness of contemporary innovation processes. Science and Public Policy, 26(1):63–64, February 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nelis:1999:WGW

[Nel99b] Annemiek Nelis. Who is governing what? Science and Public Policy, 26(5):376–378, October 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Newby:1992:OSO

[New92] Howard Newby. One society, one Wissenschaft: a 21st century vision. Science and Public Policy, 19(1):7–14, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Niosi:1991:CNS

[Nio91] Jorge Niosi. Canada’s national system of innovation. Science and Public Policy, 18(2): 83–92, April 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Noble:1993:LLP

[Nob93] Denis Noble. The logic of life: the public perception of science and its threat to the values of society. Science and Public Policy, 20(3):187–192, June 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nowotny:1994:PPS

[Now94] Helga Nowotny. Phases or paradigms of science policy? A response. Science and Public Policy, 21(6):415–416, December 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

ODoherty:1990:SAS

[O’D90] Dermot O’Doherty. Strategic alliances — an SME and small economy perspective. Science and Public Policy, 17(5):303– 310, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ogbimi:1990:PCS

[Ogb90] F. E. Ogbimi. Preparing for commercialisation of scientific research results in Nigeria. Science and Public Policy, 17(6): 373–379, December 1990. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Oakey:1990:DPI

[OHB90] R. P. Oakey, P. G. Hare, and K. Balázs. Diffusion of process innovations in an East European economy: some conceptual observations. Science and Public Policy, 17(2):97–104, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Oakey:1992:CNC

[OHB92] R. P. Oakey, P. G. Hare, and K. Balazs. Computer numerically controlled machinery diffusion within the Hungarian economy. Science and Public Policy, 19(4):241–250, August 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Oldham:1991:S

[Old91] Geoffrey Oldham. SPRU in the 1990s. Science and Public Policy, 18(6):334, December 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Packer:1994:AIR

[Pac94] Kathryn Packer. Academicindustry relations selected bibliography. Science and Public Policy, 21(2):117–119, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Palladino:1994:WS

[Pal94] Paolo Palladino. Where are the sociologists? Science and Public Policy, 21(4):279–280, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Paquet:1990:IDF

[Paq90] Gilles Paquet. Internationalization of domestic firms and governments: Anamorphosis of a palaver. Science and Public Policy, 17(5):327–332, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pavitt:1992:ITI

[Pav92] Keith Pavitt. Internationalisation of technological innovation. Science and Public Policy, 19(2): 119–123, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Phillips:1990:ISA

[Phi90] Bruce Phillips. Innovations in South Africa: a basis for technological policy guidelines. Science and Public Policy, 17(1): 27–34, February 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Piganiol:1991:LFF

[Pig91] Pierre Piganiol. Laying the foundations of French science policies. Science and Public Policy, 18(1):23–30, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Plaz:1991:DR

[Pla91] Reinaldo Plaz. Disincentives to research. Science and Public Policy, 18(5):329–330, October 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Porter:1993:ONA

[Por93] Alan L. Porter. The opening of the ‘nano-age’? Science and Public Policy, 20(5): 361–362, October 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pownall:1995:CIP

[Pow95] Ian Pownall. The capture of internalisation as a policy tool: the case of ESPRIT. Science and Public Policy, 22(1):39–49, February 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pavitt:1991:TSW

[PP91] Keith Pavitt and Pari Patel.

Technological strategies of the world’s largest companies. Science and Public Policy, 18(6): 363–368, December 1991. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Porter:1990:NSD

[PR90] Alan L. Porter and Frederick A.

Rossini. Nanotechnology: Scenarios of development and impact. Science and Public Policy, 17(4):229–234, August 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Piskunov:1992:TBS

[PS92] Dimitry I. Piskunov and Boris Saltykov. Transforming the basic structures and operating mechanisms of Soviet science. Science and Public Policy, 19(2): 111–118, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Radosevic:1991:TPT

[Rad91] Slavo Radosevic. Technology policy for the 90s: 25 tips for a policy maker. Science and Public Policy, 18(4):251–258, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1990:VBS

[Rah90] A. Rahman. Vision of a better society. Science and Public Policy, 17(1):54–56, February 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1991:SCT

[Rah91] A. Rahman. Science and the cultural transformation: a history from India. Science and Public Policy, 18(1):51–55, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rajeswari:1990:STS

[Raj90] Dr A. R. Rajeswari. Science and technology statistical system and data collection in India: Methodology, issues and problems. Science and Public Policy, 17(1):35–44, February 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ramanathan:1994:IAC

[Ram94a] K. Ramanathan. An integrated approach for the choice of appropriate technology. Science and Public Policy, 21(4):221– 233, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ramsay:1994:LIR

[Ram94b] Maureen Ramsay. Lack of intellectual rigour. Science and Public Policy, 21(2):133–134, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Randall:1994:SC

[Ran94a] D. W. Randall. Superficial and confused. Science and Public Policy, 21(5):350–351, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Randall:1994:EAP

[Ran94b] Dave Randall. Error analysis and policy decisions. Science and Public Policy, 21(4):272– 273, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rappert:1995:SNA

[Rap95a] Brian Rappert. Shifting notions of accountability in publicand private-sector research in the UK: Some central concerns. Science and Public Policy, 22(6): 383–390, December 1995. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rappert:1995:UEP

[Rap95b] Brian Rappert. The US extension of plant variety protection: a critical evaluation. Science and Public Policy, 22(2): 95–105, April 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rappert:1999:SWR

[Rap99] Brian Rappert. Science, warfare and responsibility. Science and Public Policy, 26(3):214– 215, June 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Reiss:1990:SC

[Rei90] Edward Reiss. Space for choice. Science and Public Policy, 17(4): 267–268, August 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Renn:1995:SUS

[Ren95] Ortwin Renn. Style of using scientific expertise: a comparative framework. Science and Public Policy, 22(3):147–156, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Reppy:1990:TPU

[Rep90] Judith Reppy. Thought provoking but unconvincing. Science and Public Policy, 17(5): 333–334, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Richardson:1991:QVU

[Ric91] Jacques G. Richardson. Quo vadis Unesco? An institution at the crossroads. Science and Public Policy, 18(3):142–155, June 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Richardson:1992:CME

[Ric92] Jacques G. Richardson. CERN, the mighty European laboratory for particle physics. Science and Public Policy, 19(4):251– 253, August 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Richardson:1993:IIA

[Ric93] Jacques G. Richardson. International Institute of Applied Systems Analysis, Austria. Science and Public Policy, 20(3): 211–213, June 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Richardson:1994:RCT

[Ric94a] Jacques Richardson. Research in Cuba today. Science and Public Policy, 21(3):185–187, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Richardson:1994:ROF

[Ric94b] Jacques G. Richardson. Reorienting France’s main research strategy. Science and Public Policy, 21(4):269–270, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Richardson:1994:UDW

[Ric94c] Jacques G. Richardson. Unesco and the dynamics of world science. Science and Public Policy, 21(3):197–198, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rutgers:1999:SEP

[RM99] M. R. Rutgers and M. A. Mentzel. Scientific expertise and public policy: resolving paradoxes? Science and Public Policy, 26(3):146–150, June 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Roh:1993:CVJ

[Roh93] Sung-Joon Roh. Contrasting view of Japan. Science and Public Policy, 20(1):58–59, February 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Roqueplo:1995:SEA

[Roq95] Philippe Roqueplo. Scientific expertise among political powers, administrations and public opinion. Science and Public Policy, 22(3):175–182, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rosenberg:1991:CIS

[Ros91] Nathan Rosenberg. Critical issues in science policy research. Science and Public Policy, 18(6): 335–346, December 1991. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rowlands:1994:ETP

[Row94] Mark Rowlands. Entertaining and thought-provoking. Science and Public Policy, 21(6):418– 419, December 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Reddy:1994:EPG

[RS94] A. S. Prasada Reddy and Jon Sigurdson. Emerging patterns of globalisation of corporate R&D and scope for innovative capability building in developing countries? Science and Public Policy, 21(5):283–294, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ruivo:1994:PPS

[Rui94]  Beatriz Ruivo.  ‘phases’ or

‘paradigms’ of science policy? Science and Public Policy, 21(3): 157–164, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Russell:1992:GB

[Rus92] Dr Alan Russell. Global biotechnology. Science and Public Policy, 19(2):124–125, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Russell:1995:MTP

[Rus95] Alan Russell. Merging technological paradigms and the knowledge structure in international political economy. Science and Public Policy, 22(2): 106–116, April 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ryan:1990:SAG

[Rya90] Neal Ryan. Selectivity in Australian Government support for innovation. Science and Public Policy, 17(4):210–218, August 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sagasti:1991:STP

[Sag91] Francisco R. Sagasti. Science and technology policy research: some lessons of experience and a world bank perspective. Science and Public Policy, 18(6): 379–383, December 1991. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Salomon:1995:UQM

[Sal95] Jean-Jacques Salomon. The ‘uncertain quest’: mobilising science and technology for development. Science and Public Policy, 22(1):9–18, February 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sapsed:1990:CS

[Sap90] Jon Sapsed. Check six! Science and Public Policy, 17(3): 201–202, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Saviotti:1993:EHE

[Sav93] Paolo Saviotti. Expanding the horizons of economics. Science and Public Policy, 20(4):290– 292, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stephan:1999:BDU

[SB99] Paula E. Stephan and Grant Black. Bioinformatics: Does the US system lead to missed opportunities in emerging fields? A case study. Science and Public Policy, 26(6):382–392, December 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schimank:1990:TPT

[Sch90] Uwe Schimank. Technology policy and technology transfer from state-financed research institutions to the economy: Some German experiences. Science and Public Policy, 17(4):219– 228, August 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schofield:1991:TPF

[Sch91] Steve Schofield. Towards a peaceful future. Science and Public Policy, 18(3):203–204, June 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schofield:1993:ISA

[Sch93a] Dr Steven Schofield. Impressive statistical analysis. Science and Public Policy, 20(3): 214–215, June 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schwarz:1993:TCC

[Sch93b] Dr Michiel Schwarz. The technological culture: challenges for technology assessment and policy. Science and Public Policy, 20(6):381–388, December 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Senack:1990:IR

[SdlMK90] Joy Senack, Dr John de la Mothe, and Dr Micheal Kelly. The internationalization of R&D. Science and Public Policy, 17(5):270, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Salles-Filho:1995:DBB

[SF95] Sergio Salles-Filho. Disconnections between biotechnology policies and market reality: Comments on the Brazilian experience. Science and Public Policy, 22(3):208–209, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sharp:1991:ER

[Sha91] Margaret Sharp. Europe — a renaissance? Science and Public Policy, 18(6):393–400, December 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sigurdson:1992:IRD

[Sig92] Jon Sigurdson. Internationalising research and development in Japan. Science and Public Policy, 19(3):134–144, June 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sikka:1990:FYI

[Sik90] Dr Pawan Sikka. Forty years of Indian science. Science and Public Policy, 17(1):45–53, February 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sikka:1995:SPF

[Sik95] Dr Pawan Sikka. Science policy formulation and implementation in India. Science and Public Policy, 22(6):363–368, December 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Simpson:1992:TDF

[Sim92] S. Simpson. Three-dimensional framework. Science and Public Policy, 19(4):257–259, August 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Simpson:1993:CTA

[Sim93a] Seamus Simpson. Comprehensive and timely analysis. Science and Public Policy, 20(6): 426–427, December 1993. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Simpson:1993:RTN

[Sim93b] Seamus Simpson. Rich tapestry of networks. Science and Public Policy, 20(2):133–134, April 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Simpson:1994:CMR

[Sim94a]  Seamus Simpson.  Compre-

hensive and meticulously researched. Science and Public Policy, 21(5):347–348, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Simpson:1994:EUE

[Sim94b] Seamus Simpson. Evolving and uncertain environment. Science and Public Policy, 21(2):134– 136, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Simpson:1995:GPC

[Sim95] Seamus Simpson. Government policy crucial for development. Science and Public Policy, 22(1): 69–70, February 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sinclair:1990:IAE

[Sin90] Craig Sinclair. Introduction to the Academia Europaea. Science and Public Policy, 17(2): 67–68, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sinclair:1993:ECS

[Sin93] Craig Sinclair. European cultures and science policy: The establishment of an Academy for Europe. Science and Public Policy, 20(5):343–350, October 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sirilli:1992:TBP

[Sir92] Giorgio Sirilli. Technological balance of payments as an indicator of technology transfer. Science and Public Policy, 19(6): 347–356, December 1992. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Skoie:1993:STA

[Sko93a] Hans Skoie. Science and technology advice to governments — a Norwegian perspective. Science and Public Policy, 20(2): 79–86, April 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Skolnikoff:1993:NIT

[Sko93b] Professor Eugene B. Skolnikoff.

New international trends affecting science and technology. Science and Public Policy, 20(2): 115–125, April 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Skolnikoff:1995:EUS

[Sko95] Eugene B. Skolnikoff. Evolving US science and technology policy in a changing international environment. Science and Public Policy, 22(2):74–84, April 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Slowe:1991:PVP

[Slo91] Peter Slowe. Private vs public.

Science and Public Policy, 18(5): 327–328, October 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1990:MRE

[Smi90a] Alexander McCall Smith. Medical research and ethics. Science and Public Policy, 17(2): 121–122, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1990:EMW

[Smi90b] George Teeling Smith. East meets West in medicine. Science and Public Policy, 17(4): 266–267, August 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1991:PTE

[Smi91a] Denis Smith. Power of technical expertise. Science and Public Policy, 18(2):139–140, April 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1991:MPB

[Smi91b] George Teeling Smith. Maximising potential benefits. Science and Public Policy, 18(4): 268, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1991:MS

[Smi91c] George Teeling Smith. Medicine and science. Science and Public Policy, 18(1):9–10, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1991:RIS

[Smi91d] George Teeling Smith. Rewards for innovation: Striking a balance for pharmaceuticals. Science and Public Policy, 18(1): 76–79, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smit:1993:IMT

[Smi93] Wim A. Smit. Intervening in military technological development: a comment on Donald MacKenzie’s inventing accuracy. Science and Public Policy, 20(6): 396–404, December 1993. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1994:GCC

[Smi94] Dr Thomas B. Smith. Global climate change in asia: the politics of public policy-making and science agenda setting. Science and Public Policy, 21(4): 249–259, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sanz-Menendez:1993:VSS

[SMMG93] Luis Sanz-Menéndez, Emilio Muñoz, and Clara E. García. The vicissitudes of Spanish science and technology policy: coordination and leadership. Science and Public Policy, 20(6): 370–380, December 1993. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Seidl:1999:CAT

[SP99] Peter Rudolf Seidl and Waldimir Pirro e Longo. Comments on the application of the Triple Helix of innovation to developing countries. Science and Public Policy, 26(2):137–138, April 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Spinardi:1992:DTE

[Spi92] Graham Spinardi. Defence technology enterprises: a case study in technology transfer. Science and Public Policy, 19(4):198– 206, August 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Srinivas:1999:ICI

[Sri99] K. Ravi Srinivas. Influence of conflicting interests in policymaking. Science and Public Policy, 26(6):445–446, December 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stankiewicz:1994:SCU

[Sta94] Rikard Stankiewicz. Spin-off companies from universities. Science and Public Policy, 21 (2):99–107, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stewart:1991:ESI

[Ste91] Alison C. Stewart. Ethical and social implications of the Human Genome Project: The issues in the UK. Science and Public Policy, 18(2):123–129, April 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stead:1992:CSS

[Ste92] H. Stead. Collection of S&T statistics. Science and Public Policy, 19(5):275–280, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stein:1999:ECW

[Ste99] Josephine Anne Stein. Electronic commerce: would you buy IT? Science and Public Policy, 26(1):67–68, February 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stromholm:1990:HVP

[Str90] Stig Strömholm. Hero or villain? Prometheus reconsidered. Science and Public Policy, 17 (2):69–72, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sundbo:1995:TPI

[Sun95] Jon Sundbo. Three paradigms in innovation theory. Science and Public Policy, 22(6):399– 410, December 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sanchez:1991:RDE

[SV91] Dr M. Paloma Sánchez and Dr José Vicens. Recent developments in the export of technology by Spanish companies. Science and Public Policy, 18(5): 281–293, October 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shackley:1995:GCC

[SW95a] Simon Shackley and Brian Wynne. Global climate change: the mutual construction of an emergent science-policy domain. Science and Public Policy, 22(4): 218–230, August 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shackley:1995:SSB

[SW95b] Simon Shackley and Brian Wynne. Simon Shackley and Brian Wynne’s response to the letters from Sonja Boehmer-Christiansen, and Marjolein van Asselt and Jan Rotmans. Science and Public Policy, 22(6): 415–416, December 1995. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Spinardi:1995:TDI

[SWG95] Graham Spinardi, Robin Williams, and Ian Graham. Technical data interchange in the Eurofighter project. Science and Public Policy, 22(1):29–38, February 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Szanto:1994:EAE

[Szá94] Borisz Szántó. Evolutionary aspects of East-European changes. Science and Public Policy, 21(6): 381–385, December 1994. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Taylor:1993:SFF

[Tay93] Jane Taylor. Subterranean fact and fantasy. Science and Public Policy, 20(4):289–290, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tent:1990:ICO

[Ten90] M. Tent. International cooperation in R&D: The experiences of the European Community. Science and Public Policy, 17(5):293–295, October 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Theys:1995:DME

[The95] Dr Jacques Theys. Decisionmaking on a European scale: What has changed in the relation between science, politics and expertise? Science and Public Policy, 22(3):169–174, June 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tindimubona:1993:ESC

[Tin93] Alex R. Tindimubona. Establishment of a science culture in Africa. Science and Public Policy, 20(4):245–250, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teo:1999:SII

[TL99] Thompson S. H. Teo and Vivien K. G. Lim. Singapore — an ‘intelligent island’: moving from vision to reality with information technology. Science and Public Policy, 26(1):27–36, February 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Park:1999:TNS

[tP99] Yong tae Park. A taxonomy of national systems of innovation: R&D structure of OECD economies. Science and Public Policy, 26(4):241–246, August 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Triendl:1999:NLC

[Tri99a] Robert Triendl. National laboratories in changing economic environment. Science and Public Policy, 26(6):446–447, December 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Triendl:1999:PAL

[Tri99b] Robert Triendl. Past achievements with little insight into the present. Science and Public Policy, 26(3):212–214, June 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tait:1999:PAR

[TW99] Joyce Tait and Robin Williams. Policy approaches to research and development: foresight, framework and competitiveness. Science and Public Policy, 26(2): 101–112, April 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tweedale:1990:BLW

[Twe90] Geoffrey Tweedale. Britain led the way. Science and Public Policy, 17(3):203–204, June 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Uttley:1991:BHP

[Utt91] Matthew R. H. Uttley. British helicopter production, 1945– 1960: a case study of technology transfer and market dominance. Science and Public Policy, 18(4):

235–243, August 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vanAsselt:1995:CSB

[vAR95] Marjolein B. A. van Asselt and Jan Rotmans. Comments on S. Shackley and B. Wynne paper “Global climate change: The mutual construction of an emergent sciencepolicy domain”. Science and Public Policy, 22(6):414–415, December 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vanEeten:1999:DDS

[vE99] M. J. G. van Eeten. ‘dialogues of the deaf’ on science in policy controversies. Science and Public Policy, 26(3):185–192, June 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vernon:1994:BSK

[Ver94] Keith Vernon. Bringing scientific knowledge to the masses. Science and Public Policy, 21(3): 195–196, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vernon:1995:WHT

[Ver95] Keith Vernon. What happened to technical education? Science and Public Policy, 22(5): 350–351, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vanHulst:1995:DES

[vHdP95] Noé van Hulst and Koen de Pater. Does Europe still matter in the technology race? Science and Public Policy, 22(2): 117–120, April 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Visalakshi:1995:RCS

[Vis95] S. Visalakshi. R&D and commercialisation status of hybridoma technology in India: an international comparison. Science and Public Policy, 22(4): 231–238, August 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vitta:1992:URS

[Vit92] Paul B. Vitta. Utility of research in sub-Saharan Africa: Beyond the leap of faith. Science and Public Policy, 19(4): 221–228, August 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vonortas:1995:NDU

[Von95] Nicholas S. Vonortas. New directions for US science and technology policy: the view from the R&D assessment front. Science and Public Policy, 22(1):19–28, February 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Voyer:1995:GRI

[Voy95a]  Roger Voyer.  Globalization

and regional industrial clustering. Science and Public Policy, 22(5):347–348, October 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Voyer:1995:RGR

[Voy95b] Roger Voyer. Rugby game not relay race. Science and Public Policy, 22(6):418–419, December 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Voyer:1999:TYC

[Voy99] Roger Voyer. Thirty years of Canadian science policy: from 1.5 to 1.5. Science and Public Policy, 26(4):277–282, August 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Walker:1991:TNW

[Wal91] William Walker. The Triad and the New World Order. Science and Public Policy, 18(6): 401–406, December 1991. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Walshe:1992:RDT

[Wal92] Grahame Walshe. Research and development trends: criteria for assessment. Science and Public Policy, 19(2):75–88, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Walsh:1993:DPM

[Wal93] Vivien Walsh. Demand, public markets and innovation in biotechnology. Science and Public Policy, 20(3):138–156, June 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watson:1994:ALD

[Wat94a] Adrian Watson. Ambitious but lacking depth. Science and Public Policy, 21(3):194–195, June 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watson:1994:CSI

[Wat94b] Adrian Watson. Can science influence environmental policy? Science and Public Policy, 21(5): 351–352, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Webster:1990:ISE

[Web90] A. J. Webster. Institutional stability: Engineering an environment for biotechnology. Science and Public Policy, 17(6): 381–386, December 1990. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Webster:1994:BIR

[Web94a] Andrew Webster. Bridging institutions: the role of contract research organisations in technology transfer. Science and Public Policy, 21(2):89–97, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Webster:1994:IEA

[Web94b] Andrew Webster. International evaluation of academic-industry relations: contexts and analysis. Science and Public Policy, 21 (2):72–78, April 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Weingart:1999:SEP

[Wei99] Peter Weingart. Scientific expertise and political accountability: paradoxes of science in politics. Science and Public Policy, 26(3): 151–161, June 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Welsh:1994:LRT

[Wel94] Ian Welsh. Letting the research tail wag the end-user’s dog: The Powell Committee and UK nuclear technology. Science and Public Policy, 21(1):43–53, February 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Weyer:1994:ANH

[Wey94] Johannes Weyer. Actor networks and high risk technologies: the case of the Gulf War. Science and Public Policy, 21(5): 321–334, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wakabayashi:1999:SRA

[WGBW99] Koji Wakabayashi, Charla Griffy-Brown, and Chihiro Watanabe. Stimulating R&D: an analysis of the Ministry of International Trade and Industry’s ‘visions’ and the current challenges facing Japan’s technology policy-making mechanisms. Science and Public Policy, 26(1): 2–16, February 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Whyte:1994:CWI

[Why94] Neil Whyte. Competent weaving of information. Science and Public Policy, 21(6):424–425, December 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilson:1990:SCP

[Wil90] Andrew H. Wilson. Science culture and public policy: myths, realities and prospects. Science and Public Policy, 17(2): 111–120, April 1990. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Williams:1992:RSW

[Wil92] Roger Williams. R&D statistics: the work of the House of Lords. Science and Public Policy, 19(5): 291–295, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilson:1993:UR

[Wil93] Alan Wilson. From untouchable to respected. Science and Public Policy, 20(6):427–428, December 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wheale:1992:EGE

[WM92] Peter R. Wheale and Ruth M. McNally. Ethics and genetic engineering. Science and Public Policy, 19(1):64–66, February 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wheale:1993:BPE

[WM93] Peter Wheale and Ruth Mc-Nally. Biotechnology policy in Europe: a critical evaluation. Science and Public Policy, 20(4): 261–279, August 1993. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Woherem:1991:ESM

[Woh91] Dr Evans E. Woherem. Expert systems as a medium for knowledge transfer to less developed countries. Science and Public Policy, 18(5):301–309, October 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Woolgar:1991:BCD

[Woo91] Steve Woolgar. Beyond the citation debate: towards a sociology of measurement technologies and their use in science policy. Science and Public Policy, 18(5): 319–326, October 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wang:1999:EPG

[WY99] Ming Yeu Wang and Benjamin J. C. Yuan. Evaluation practice of the government-sponsored programme for promoting industrial upgrading in Taiwan. Science and Public Policy, 26(2): 125–135, April 1999. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Yearley:1994:PCS

[Yea94] Steven Yearley. ‘post-constructivist’ science studies. Science and Public Policy, 21(5):348–350, October 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Young:1991:EPC

[You91] Stephen Young. Energy politics: 20th century controversies. Science and Public Policy, 18(6): 422–424, December 1991. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Young:1994:OFC

[You94] Craig Young. Is organic farming cleaner? Science and Public Policy, 21(4):271–272, August 1994. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Yuhai:1992:SSC

[Yuh92] Zhao Yuhai. S&T statistics collection and the role of S&T indicators in the formulation of S&T policies in China. Science and Public Policy, 19(5): 296–302, October 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zacher:1992:MPG

[Zac92] Hans F. Zacher. The Max-Planck-Gesellschaft and German unification. Science and Public Policy, 19(2):71–74, April 1992. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zaparoli:1995:PS

[Zap95] Edson Luiz Zaparoli. Power in space. Science and Public Policy, 22(4):276–277, August 1995. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zhongliang:1991:PSP

[Zho91] Zhang Zhongliang. People and science: public attitudes in China toward science and technology. Science and Public Policy, 18(5):311–317, October 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ziman:1991:NNM

[Zim91] John Ziman. A neural net model of innovation. Science and Public Policy, 18(1):65–75, February 1991. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).